### 18 Mar 2026 v2

# Positivity is undecidable in tensor products of free algebras

Abstract.
It is well known that an element of the algebra of noncommutative *-polynomials is positive in all *-representations if and only if it is a sum of squares. This provides an effective way to determine if a given *-polynomial is positive, by searching through sums of squares decompositions. We show that no such procedure exists for the tensor product of two noncommutative *-polynomial algebras: determining whether a *-polynomial of such an algebra is positive is coRE-hard. We also show that it is coRE-hard to determine whether a noncommutative *-polynomial is trace-positive. Our results hold if noncommutative *-polynomial algebras are replaced by other sufficiently free algebras such as group algebras of free groups or free products of cyclic groups.
1. Introduction
Recall that a self-adjoint bounded operator $T$ on a Hilbert space $H$ is positive if $\langle v,Tv\rangle\geq 0$ for all $v\in H$. A self-adjoint element $\alpha$ of a complex *-algebra $\mathcal{A}$ is said to be positive if $\pi(\alpha)$ is positive for all *-representations $\pi:\mathcal{A}\to\mathcal{B}(H)$. Suppose that $\mathcal{A}$ is the complexification of a finitely generated *-algebra $\mathcal{A}_{\mathbb{Q}}$ over $\mathbb{Q}$. Elements of $\mathcal{A}_{\mathbb{Q}}$ can be represented as noncommutative *-polynomials with rational coefficients, and hence we can work with them concretely, for instance on a computer.
In many situations, we’d like to be able to decide whether an element $a\in\mathcal{A}_{\mathbb{Q}}$ is positive in $\mathcal{A}$, and if it is, find a certificate of positivity. A prototypical example is $\mathcal{A}=\mathbb{C}\mathcal{F}_{n}$, the group algebra of the free group $\mathcal{F}_{n}$, with $\mathcal{A}_{\mathbb{Q}}=\mathbb{Q}\mathcal{F}_{n}$. In this case, the following two properties hold:
  1. (a)
    Every positive element $\alpha\in\mathcal{A}$ is a sum of (hermitian) squares (SOS), meaning that there are $b_{1},\ldots,b_{k}\in\mathcal{A}$ such that $\alpha=\sum_{i=1}^{k}b_{i}^{*}b_{i}$.
  2. (b)
    A self-adjoint element $\alpha\in\mathcal{A}$ is positive if and only if $\pi(\alpha)$ is positive in all finite-dimensional representations $\pi:\mathcal{A}\to\mathcal{B}(H)$.
Property (a) is called the SOS property; for archimedean algebras (see the next section for a definition) such as $\mathbb{C}\mathcal{F}_{n}$, this property is also called archimedean closed. Since a sum of squares is always positive, a sum of squares decomposition for an element $a$ can be used as a certificate of positivity. The SOS property states that every positive element has such a certificate. For archimedean algebras, property (b) states that the enveloping $C^{*}$-algebra of $\mathcal{A}$ is residually finite-dimensional (RFD), or equivalently that the quadratic module generated by hermitian squares in $\mathcal{A}$ is RFD [ANT19]. For convenience, we’ll shorten this and just say that $\mathcal{A}$ is RFD. [^1] But note that this is not equivalent to the standard definition of RFD for *-algebras, which says that all non-zero elements are non-zero in some finite-dimensional representation. If $\mathcal{A}$ is RFD, then for every non-positive self-adjoint element $\alpha\in\mathcal{A}$, there is a finite-dimensional representation $\phi$ such that $\phi(\alpha)$ is not positive. Thus if $\mathcal{A}_{\mathbb{Q}}$ is finitely presented, then this representation can be used as a certificate of non-positivity. If $\mathcal{A}$ has the SOS property and is RFD, and $\mathcal{A}_{\mathbb{Q}}$ is finitely presented, it is possible to decide whether a self-adjoint element $\alpha\in\mathcal{A}_{\mathbb{Q}}$ is positive by simultaneously searching for a sum of squares decomposition for $\alpha$, and for a finite-dimensional representation showing that $\alpha$ is not positive.
For the free group algebra, property (a) is sometimes attributed to an unpublished result of Schmüdgen from the 1980s (see [NT13, OZA13]), and can also be found in [MCC01, HMP04, BT07, NT13, OZA13]. Property (b) is a result of Choi [CHO80]. Both properties also hold for the *-algebra of noncommutative *-polynomials $\mathbb{C}^{*}\langle x_{1},\ldots,x_{n}\rangle$ by theorems of Helton [HEL02] and McCullough [MCC01], and for the *-algebra of $n$ contractions by theorems of Helton-McCullough-Putinar [HMP04] or Helton-Klep-McCullough [HKM12]. For the group algebra of a virtually free group such as the free product $\mathbb{Z}_{m}^{*n}$ of cyclic groups, a proof of both properties is again attributed to Schmüdgen [NT13], with a complete proof appearing in [COH23]. That all these algebras are RFD also follows from [GM90] or [EL92]. Thus it is possible to decide positivity for all these algebras using the algorithm above, although actually we can do better: the proofs of the SOS property above are effective, meaning that they provide bounds on the degree of polynomials in the sum of squares needed to show that an element is positive. Hence we can decide positivity by searching just through SOS decompositions and stopping when the bound is reached.
In comparison, the commutative polynomial rings $\mathbb{C}[x_{1},\ldots,x_{n}]$ famously do not have the SOS property for $n\geq 2$. However, Artin’s solution of Hilbert’s 17th problem states that every positive polynomial is a sum of squares of rational functions [ART27]. More generally, a result of Scheiderer [SCH00] states that finitely generated connected commutative *-algebras of dimension $\geq 3$, such as the group algebra $\mathbb{C}\mathbb{Z}^{n}$ for $n\geq 3$, do not have the SOS property (provided they have a state). Certificates of positivity are provided for all finitely generated commutative algebras by the Krivine-Stengle Positivstellensatz [KRI64, STE74]. Effective versions of Artin’s theorem and the Krivine-Stengle Positivstellensatz are proven in [LPR20], and although the degree bounds are too large to yield practical algorithms, this gives a way to decide positivity in principle. For finitely generated commutative algebras, positivity can also be decided with Tarski quantifier elimination or other algorithms for the first-order theory of the reals. In fact, since non-positivity lies in the existential theory of the reals, positivity can be decided in PSPACE [CAN88].
If $\mathcal{A}$ is a finitely generated commutative archimedean *-algebra, then Schmüdgen’s Positivstellensatz states that every element which is strictly positive on the state space of $\mathcal{A}$ is a sum of squares [SCH91]. For all archimedean *-algebras $\mathcal{A}$, Helton and McCullough’s Positivstellensatz [HM04] says that $\alpha\in\mathcal{A}$ is positive if and only if $\alpha+\epsilon$ is a sum of squares for all $\epsilon>0$ (this exact formulation seems to appear first in [CIM09], see also [SCH09, OZA13]). This does not give a method for deciding positivity, but does show that deciding positivity is in the class $\Pi_{2}^{0}$ for finitely presented archimedean algebras.
While positivity is decidable for the free group algebras and other prototypical noncommutative *-algebras $\mathcal{A}$ listed above, the point of this paper is to show that positivity becomes undecidable for tensor products $\mathcal{A}\otimes\mathcal{A}$ of such algebras:
Theorem 1.1.
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}=\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathbb{C}$ be the complexification. Then the decision problem
(Pos) Given $\alpha\in\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$, is $\alpha$ positive as an element of $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$?
is coRE-hard.
Here, $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ denotes the *-algebra of noncommutative *-polynomials over $\mathbb{Q}$. Specifically, Theorem 1.1 shows that there is a computable mapping from Turing machines $M$ to self-adjoint elements $\alpha_{M}\in\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$, such that $\alpha_{M}$ is positive if and only if $M$ does not halt. The bounds on $n$ and $m$ in Theorem 1.1 are tight, as the algebras $\mathbb{Q}^{*}\langle x\rangle\otimes\mathbb{Q}^{*}\langle x\rangle=\mathbb{Q}[x_{1},x_{2}]$, $\mathbb{Q}\mathcal{F}_{1}\otimes\mathbb{Q}\mathcal{F}_{1}=\mathbb{Q}\mathbb{Z}\times\mathbb{Z}$, and $\mathbb{Q}\mathbb{Z}_{m}\times\mathbb{Q}\mathbb{Z}_{m}$ are commutative, and positivity is also decidable for $\mathbb{Q}\mathbb{Z}_{2}^{*2}\times\mathbb{Z}_{2}^{*2}$ (see Example 6.8). As an immediate corollary of Theorem 1.1, we get:
Corollary 1.2.
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}$ be the complexification. Then there are elements of $\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$ which are positive, but are not sums of hermitian squares in $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$, so $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$ does not have the SOS property.
Proof.
If every element of $\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$ is a sum of squares in $\mathcal{A}\otimes_{\mathbb{C}}\mathcal{A}$, then the problem of deciding whether elements of $\mathcal{A}\otimes\mathcal{A}$ are positive is contained in RE by searching through SOS decompositions. Since coRE is not contained in RE, this is not possible. ∎
For the algebras $\mathbb{Q}\mathcal{F}_{n}$ and $\mathbb{Q}\mathbb{Z}_{m}^{*n}$, $m\geq 3$, the hypothesis $n\geq 2$ in 1.2 is necessary, since finite-dimensional $C^{*}$-algebras like $\mathbb{C}\mathbb{Z}_{m}\times\mathbb{Z}_{m}$ always have the SOS property, and $\mathbb{C}\mathbb{Z}\times\mathbb{Z}$ has the SOS property by a result of Scheiderer [SCH06]. We do not know if $\mathbb{C}\mathbb{Z}_{2}^{*2}\times\mathbb{Z}_{2}^{*2}$ has the SOS property. For $\mathcal{A}_{\mathbb{Q}}=\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$, the corollary already follows (for all $n\geq 1$) from the fact that $\mathbb{C}[x,y]$ does not have the SOS property; indeed, if $p(x,y)$ is a positive polynomial in $\mathbb{C}[x,y]$ which is not a sum of squares, then $p(x_{1}\otimes 1,1\otimes x_{1})$ is positive in $\mathcal{A}\otimes\mathcal{A}$, but does not have a sum of squares decomposition. The Motzkin polynomials provide explicit examples of positive polynomials in $\mathbb{Q}[x,y]$ which are not sums of squares, although of course they are sums of squares of rational functions by Artin’s theorem. 1.2 does not seem to have been known for $\mathcal{A}_{\mathbb{Q}}=\mathbb{Q}\mathcal{F}_{n}$ and $\mathbb{Q}\mathbb{Z}_{m}^{*n}$; however, Theorem 1.1 is stronger, in that SOS decompositions can be replaced in 1.2 with any other certificate of positivity. For instance, for $\mathcal{A}_{\mathbb{Q}}=\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ with $n\geq 2$, there are positive elements of $\mathcal{A}_{\mathbb{Q}}\otimes_{\mathbb{Q}}\mathcal{A}_{\mathbb{Q}}$ which are not sums of squares of noncommutative rational functions in the sense of [KSV23]. The proof of Theorem 1.1 is given in Section 6. We actually prove a more general statement (Theorem 6.1), which applies to other algebras including the algebra of bounded contractions.
The proof of Theorem 6.1 is by reduction to tracial positivity. An element $\alpha$ of a *-algebra $\mathcal{A}$ is said to be trace-positive if $\tau(\alpha)\geq 0$ for all (bounded) tracial states $\tau$ on $\mathcal{A}$. We show that tracial positivity is undecidable for free *-polynomials and group algebras:
Theorem 1.3.
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{m}^{*N}$ for $m\geq 2$. If $N$ is large enough, then the decision problem
(TrPos) Given $\alpha\in\mathcal{A}_{\mathbb{Q}}$, is $\alpha$ trace-positive as an element of $\mathcal{A}$?
is coRE-hard.
Klep and Schweighofer [KS08] prove an analog of Helton and McCullough’s Positivstellensatz for the algebra of bounded contractions: an element $\alpha$ is trace-positive if and only if $\alpha+\epsilon$ is a sum of hermitian squares and commutators for all $\epsilon>0$. This is extended to free group algebras by Juschenko and Popovych [JP11], to archimedean algebras by Ozawa [OZA13], and to $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ by Klep, Scheiderer, and Volčič [KSV23]. Quarez [QUA15] (see also [KSV23, Example 3.2]) gives examples of trace-positive polynomials in $\mathbb{Q}^{*}\langle x_{1},x_{2}\rangle$ which are not sums of hermitian squares and commutators, but which are sums of squares and commutators of noncommutative rational functions. As in 1.2, Theorem 1.3 immediately implies that all the algebras listed in the theorem contain trace-positive elements which are not sums of squares and commutators of noncommutative rational functions. Unlike Theorem 1.1, we do not know the smallest bound $N$ such that Theorem 1.3 holds.
Part of the motivation for Theorem 1.1 comes from Kirchberg’s reformulation of the Connes embedding problem, sometimes called Kirchberg’s conjecture [KIR93], and the recent MIP*=RE theorem of Ji, Natarajan, Vidick, Wright, and Yuen [JNV+20], which resolves this conjecture. Kirchberg’s conjecture asks whether the max and min tensor norms on $\mathbb{C}\mathcal{F}_{n}\otimes\mathcal{F}_{n}=\mathbb{C}\mathcal{F}_{n}\times\mathcal{F}_{n}$ agree for $n\geq 2$ [KIR93]. As Ozawa observes [OZA04, OZA13], if $\mathcal{A}$ is archimedean and RFD, then the max and min tensor norms on $\mathcal{A}\otimes\mathcal{A}$ agree if and only if $\mathcal{A}\otimes\mathcal{A}$ is RFD, so Kirchberg’s conjecture is equivalent to whether $\mathbb{C}\mathcal{F}_{n}\otimes\mathbb{C}\mathcal{F}_{n}=\mathbb{C}\mathcal{F}_{n}\times\mathcal{F}_{n}$ is RFD. Kirchberg’s conjecture is still equivalent to the Connes embedding problem if $\mathcal{F}_{n}$ is replaced by $\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$ [OZA13, FRI12]. The algebras $\mathbb{C}\mathbb{Z}_{m}^{*n}\otimes\mathbb{C}\mathbb{Z}_{m}^{*n}$ are of particular interest in quantum information, since they correspond to bipartite (meaning composed of two spatially separated subsystems) measurement scenarios with $n$ measurement settings and $m$ measurement outcomes per subsystem, and they are closely connected with the complexity class MIP* of multiprover interactive proof systems with entangled provers.
The main result of [JNV+20] is that MIP* is equal to RE, the class of recursively enumerable languages. In algebraic language, this implies that there is a computable mapping from Turing machines $M$ to elements $\alpha_{M}\in\overline{\mathbb{Q}}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ (where $\overline{\mathbb{Q}}$ is the algebraic closure of $\mathbb{Q}$, and $m$ and $n$ also depend on $M$), such that if $M$ halts then $\alpha_{M}$ is not positive in finite-dimensional representations of $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$, and if $M$ does not halt, then $\alpha_{M}-1$ is positive in finite-dimensional representations of $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$. This shows that the promise problem
(GapFDPos) Given $n,m\geq 1$ and $\alpha\in\overline{\mathbb{Q}}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$, determine if either $\alpha-1$ is positive in finite-dimensional representations, or $\alpha$ is not positive in some finite-dimensional representation, promised that one of the two is the case.
is coRE-hard, so it is undecidable to determine if an element of $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ is strictly positive in finite-dimensional representations. If $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ were RFD, then (GapFDPos) would be equivalent to the promise problem
(GapPos) Given $n,m\geq 1$ and $\alpha\in\overline{\mathbb{Q}}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$, determine if either $\alpha-1$ is positive, or $\alpha$ is not positive, promised that one of the two is the case.
Note that $\alpha-1$ is positive if and only if $\alpha$ is strictly positive with spectrum contained in $1,+\infty)$, so (GapPos) is a “gapped” version of the decision problem (Pos) considered in Theorem 1.1 (albeit with $\mathbb{Q}$ replaced by $\overline{\mathbb{Q}}$, and $n$ and $m$ allowed to vary). By Helton and McCullough’s Positivstellensatz, if $\alpha-1$ is positive then $\alpha$ has an SOS decomposition proving positivity. Due to the promise in (GapPos), a proof that $\alpha$ is positive is also a proof that $\alpha-1$ is positive. So unlike (Pos), the problem (GapPos) is contained in RE. Since RE does not contain coRE, (GapPos) cannot be coRE-hard, and hence the MIP*=RE theorem implies that $\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ is not RFD for large enough $n$ and $m$.
There is a natural complement to the $\text{MIP}^{*}=\text{RE}$ theorem called the $\text{MIP}^{\text{co}}=\text{coRE}$ conjecture, which if true would imply (GapPos) is RE-complete, and hence that (Pos) is RE-hard for $\mathcal{A}=\mathbb{C}\mathbb{Z}_{m}^{*n}$ (assuming again that $\mathbb{Q}$ is replaced by $\overline{\mathbb{Q}}$, and $n$ and $m$ are allowed to vary). If $\mathcal{A}\otimes\mathcal{A}$ is RFD, then (Pos) is contained in coRE, so showing that (Pos) is RE-hard for $\mathcal{A}=\mathbb{C}\mathbb{Z}_{m}^{*n}$ (with or without varying $n$ and $m$) would give another disproof of the Connes embedding problem. Proving that (GapPos) is RE-complete would also imply that it is undecidable to determine if $\alpha\in\mathbb{C}\mathbb{Z}_{m}^{*n}\times\mathbb{Z}_{m}^{*n}$ is a sum of squares. On the other hand, because (GapPos) is contained in RE, hardness results for (GapPos) cannot be used to prove 1.2, and Theorem 1.1 is not true if (Pos) is replaced by (GapPos). Hence Theorem 1.1 and the $\text{MIP}^{\text{co}}=\text{coRE}$ conjecture seem to be complementary. In any case, the proof methods we used here are different from the methods used in the proof of $\text{MIP}^{*}=\text{RE}$, and our techniques do not appear to be applicable to the $\text{MIP}^{\text{co}}=\text{coRE}$ conjecture.
As mentioned above, the only upper bound we have on (Pos) is $\Pi^{0}_{2}$. If (Pos) contains both RE and coRE, then it seems possible that it is equal to $\Pi^{0}_{2}$, and we conjecture that this is the case for all the algebras in Theorem 1.1.
Conjecture 1.4.
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n}\rangle$ or $\mathbb{Q}\mathcal{F}_{n}$ for $n\geq 2$, or $\mathbb{Q}\mathbb{Z}_{m}^{*n}$ for $n\geq 3$, $m\geq 2$ or $n\geq 2$, $m\geq 3$, and let $\mathcal{A}$ be the complexification. Then the decision problem (Pos) from Theorem 1.1 is $\Pi_{2}^{0}$-complete.
Tracial positivity has also been studied in connection with the Connes embedding problem [HAD01, KS08, BDK+14, JP11, OZA13], and the above connection with $\text{MIP}^{*}=\text{RE}$ and $\text{MIP}^{\text{co}}=\text{coRE}$ could be described equally as well in terms of tracial positivity and MIP protocols with synchronous strategies (see e.g. [HMP+19, KPS18, MNY22, NZ23]). For tracial positivity, we make the following conjecture.
Conjecture 1.5.
Let $\mathcal{A}_{\mathbb{Q}}$ be one of the algebras $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{m}^{*N}$ for $m\geq 2$, and let $\mathcal{A}$ be the complexification. If $N$ is large enough, then the decision problem (TrPos) from Theorem 1.3 is $\Pi_{2}^{0}$-complete.
If 1.5 is true, an additional problem is to find, for each algebra listed above, the smallest bound on $N$ such that the decision problem (TrPos) is $\Pi_{2}^{0}$-complete.
1.1. Acknowledgements
We thank Igor Klep, Tim Netzer, Andreas Thom, and Jurij Volčič for helpful conversations. We thank the anonymous referee for helpful comments, including supplying reference [SS11]. AM is supported by NSERC Alliance Consortia Quantum grants, reference number: ALLRP 578455 - 22. WS is supported by NSERC DG 2018-03968 and an Alfred P. Sloan Research Fellowship.
2. Notation and background
2.1. Algebras and states
We work with both groups and *-algebras. If $\mathbb{K}$ is a field with an involution, then a *-algebra over $\mathbb{K}$ is a unital associative $\mathbb{K}$-algebra $\mathcal{A}$ with an antilinear involution $\mathcal{A}\rightarrow\mathcal{A}:a\mapsto a^{*}$ such that $(ab)^{*}=b^{*}a^{*}$ for all $a,b\in\mathcal{A}$. A *-homomorphism $\pi:\mathcal{A}\to\mathcal{B}$ between *-algebras $\mathcal{A}$ and $\mathcal{B}$ is an algebra homomorphism such that $\pi(a^{*})=\pi(a)^{*}$. In this paper, $\mathbb{K}$ will be either $\mathbb{C}$ or $\mathbb{Q}$ (with trivial involution), with $\mathbb{C}$ being the default choice. When $\mathcal{A}$ is a *-algebra over $\mathbb{Q}$, we use $\mathcal{A}_{\mathbb{C}}$ to denote $\mathcal{A}\otimes_{\mathbb{Q}}\mathbb{C}$. We use $1$ for the identity in both groups and *-algebras.
If $G$ is a group, then we let $\mathbb{K}G$ be the group algebra of $G$ over the field $\mathbb{K}$. Recall that this is a *-algebra with $g^{*}=g^{-1}$. If $S$ is a set, then $\mathcal{F}_{S}$ will denote the free group with generating set $S$, and $\mathcal{F}_{n}:=\mathcal{F}_{\{x_{1},\ldots,x_{n}\}}$. We use the standard notation for group presentations, in which if $S$ is a set, and $R\subseteq\mathcal{F}_{S}$, then $\langle S:R\rangle$ denotes the quotient of $\mathcal{F}_{S}$ by the normal subgroup generated by $R$. We often want to add generators and relations to a given presentation $G=\langle S:R\rangle$, so if $S_{1}$ is another set disjoint from $S$, and $R_{1}\subseteq\mathcal{F}_{S\cup S_{1}}$, then we define $\langle G,S_{1}:R_{1}\rangle:=\langle S\cup S_{1}:R\cup R_{1}\rangle$. If $G$ is a group, then $G^{*n}$ denotes the free product of $n$ copies of $G$. The cyclic group of order $k$ will be denoted by $\mathbb{Z}_{k}$. If $S$ is a set, then $\mathbb{Z}_{k}^{*S}$ denotes the free product of $|S|$ copies of $\mathbb{Z}_{k}$, where the generators of the factors are denoted by the elements of $S$. In other words, $\mathbb{Z}_{k}^{*S}=\langle S:s^{k}=1\text{ for all }s\in S\rangle$.
We also define *-algebras by means of presentations. If $S$ is a set, we let $\mathbb{K}^{*}\langle S\rangle$ denote the free noncommutative *-algebra with generating set $S$ over $\mathbb{K}$. Elements of $\mathbb{K}^{*}\langle S\rangle$ are noncommutative *-polynomials over $\mathbb{K}$. For instance, if $S=\{x,y\}$, then elements of $\mathbb{K}^{*}\langle S\rangle$ are noncommutative polynomials in $x,x^{*},y,y^{*}$, and the involution $a\mapsto a^{*}$ exchanges $x$ with $x^{*}$ and $y$ with $y^{*}$. If $R\subseteq\mathbb{K}^{*}\langle S\rangle$, then $\mathbb{K}^{*}\langle S:R\rangle$ denotes the quotient of $\mathbb{K}^{*}\langle S\rangle$ by the two-sided *-ideal generated by $R$. As with groups, if $\mathcal{A}=\mathbb{K}^{*}\langle S:R\rangle$ is a *-algebra presentation, $S_{1}$ is a set disjoint from $S$, and $R_{1}\subseteq\mathbb{K}^{*}\langle S\cup S_{1}\rangle$, then
$\mathbb{K}^{*}\langle\mathcal{A},S:R_{1}\rangle:=\mathbb{K}^{*}\langle S\cup S_{1}:R\cup R_{1}\rangle.$
If $G=\langle S:R\rangle$ is a group presentation, then
$\mathbb{K}G=\mathbb{K}^{*}\langle S:1-r,r\in R,1-ss^{*},1-s^{*}s,s\in S\rangle.$
For both group and *-algebra presentations, we’ll follow the standard practice of writing $a=b$ for the relation $b^{-1}a$ (in the case of groups) or aba-b (in the case of *-algebras). Also, a presented group (resp. algebra) is a group (resp. algebra) with a fixed presentation, and a finitely presented group (resp. algebra) is a presented group (resp. algebra) where the presentation has a finite number of generators and relations.
As mentioned above, unless specified otherwise a *-algebra will mean a *-algebra over $\mathbb{C}$. If $\mathcal{H}$ is a Hilbert space, then $\mathcal{B}(\mathcal{H})$ is the $C^{*}$-algebra of bounded operators on $\mathcal{H}$. A representation of a *-algebra $\mathcal{A}$ is a *-homomorphism $\mathcal{A}\to\mathcal{B}(\mathcal{H})$ for some Hilbert space $\mathcal{H}$. A representation is said to be finite-dimensional if $\mathcal{H}$ is finite-dimensional (we mention this notion in the preliminary but we won’t work with it in this paper). As mentioned in the introduction, a self-adjoint element $a\in\mathcal{A}$ is (representation) positive if and only if $\phi(a)$ is positive in $\mathcal{B}(\mathcal{H})$ for all Hilbert spaces $\mathcal{H}$ and representations $\phi:\mathcal{A}\to\mathcal{B}(\mathcal{H})$.
A (bounded) state on a *-algebra $\mathcal{A}$ is a linear function $\psi:\mathcal{A}\rightarrow\mathbb{C}$ with $\psi(1)=1$ such that $\psi(a^{*})=\overline{\psi(a)}$, $\psi(a^{*}a)\geq 0$, and
(2.1) $\sup\left\{\frac{\psi(b^{*}a^{*}ab)}{\psi(b^{*}b)}:b\in\mathcal{A},\psi(b^{*}b)\neq 0\right\}<\infty$
for all $a\in\mathcal{A}$. The GNS theorem for *-algebras states that $\psi$ is a state if and only if there is a representation $\pi:\mathcal{A}\rightarrow\mathcal{B}(\mathcal{H})$ and a unit vector $v\in\mathcal{H}$ such that $\psi(a)=\langle v,\pi(a)v\rangle$ for all $a\in\mathcal{A}$ [SCH20, Theorem 4.38]. Given a state $\psi$, there is always such a representation $\pi$ and vector $v$ such that $\pi(\mathcal{A})v$ is dense in $\mathcal{H}$. A pair $(\pi,v)$ such that $\pi(\mathcal{A})v$ is dense in $\mathcal{H}$ and $\langle v,\pi(a)v\rangle=\psi(a)$ for all $a\in\mathcal{A}$ is called a GNS representation of $\psi$. The GNS representation theorem implies that a self-adjoint element $a\in\mathcal{A}$ is positive if and only if $f(a)\geq 0$ for all states $f$ on $\mathcal{A}$.
A state $\psi$ on a *-algebra $\mathcal{A}$ induces a semi-norm $\lVert a\rVert_{\psi}:=\sqrt{\psi(a^{*}a)}$ on $\mathcal{A}$ satisfying $\lvert\psi(ab)\rvert\leq\lVert a\rVert_{\psi}\lVert b\rVert_{\psi}$ for all $a,b\in\mathcal{A}$. A state $\psi$ on $\mathcal{A}$ is tracial if $\psi(ab)=\psi(ba)$ for all $a,b\in\mathcal{A}$. If $\psi$ is tracial, then $\lVert\cdotp\rVert_{\psi}$ is unitarily invariant, meaning that $\lVert uav\rVert_{\psi}=\lVert a\rVert_{\psi}$ for all unitaries $u,v\in\mathcal{A}$ and elements $a\in\mathcal{A}$. For an arbitrary state $\psi$, the seminorm $\lVert\cdotp\rVert_{\psi}$ is left unitarily invariant, but not necessarily right unitarily invariant.
An element $a$ of a *-algebra $\mathcal{A}$ is a hermitian square if there is $b\in\mathcal{A}$ such that $a=b^{*}b$, and a sum of (hermitian) squares if there is $b_{1},\ldots,b_{k}\in\mathcal{A}$ such that $a=\sum_{i}b_{i}^{*}b_{i}$. If $a$ is a sum of squares, then $\psi(a)\geq 0$ for all states $\psi$. If $a_{1}$ and $a_{2}$ are sums of squares, then $\psi(a_{1}a_{2})\geq 0$ for all tracial states $\psi$, since $\psi(b_{1}^{*}b_{1}b_{2}^{*}b_{2})=\psi((b_{2}b_{1}^{*})^{*}(b_{2}b_{1}^{*}))$ when $\psi$ is tracial.
[…]
A *-algebra $\mathcal{A}$ is said to be archimedean if for all $a\in\mathcal{A}$ (or equivalently, all $a$ in a generating set for $\mathcal{A}$), there is a scalar $\lambda>0$ such that $\lambda-a^{*}a$ is a sum of squares. If $\mathcal{A}$ is archimedean, then a linear functional $f:\mathcal{A}\to\mathbb{C}$ satisfying $f(1)=1$ and $f(a^{*})=\overline{f(a)}$, $f(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$ automatically satisfies the boundedness condition in Equation (2.1). Also, the operator norm $\lVert\cdotp\rVert_{\mathcal{A}}$ takes only finite values, and (as mentioned in the introduction) Helton and McCullough’s Positivstellensatz states that $a\in\mathcal{A}$ is positive if and only if $a+\epsilon$ is a sum of squares for all $\epsilon>0$. Group algebras $\mathbb{C}G$ are archimedean, but free *-algebras $\mathbb{C}^{*}\langle S\rangle$ are not when $S$ is non-empty, so not all the algebras we deal with are archimedean. More background can be found in [SCH09] and [OZA13].
2.2. Computability
If $\Sigma$ is a set, a word (also called a string) of length $n$ over $\Sigma$ is a finite sequence $a_{1}a_{2}\cdots a_{n}$, where $a_{i}\in\Sigma$ for all $1\leq i\leq n$. We use the standard terminology from the theory of computation (see, e.g. [SIP12]). In particular, a subset $\mathcal{L}$ of words over a finite set $\Sigma$ is in RE if there is a Turing machine $M$ with input alphabet $\Sigma$ which recognizes $\mathcal{L}$, meaning that there is a computation path where $M$ halts on input $x$ if and only if $x\in\mathcal{L}$. In this definition $M$ can be non-deterministic, but every language in RE can be recognized by a deterministic Turing machine. A language $\mathcal{L}$ is in coRE if its complement $\overline{\mathcal{L}}$ is in RE, and is decidable if it is in both RE and coRE. Equivalently, $\mathcal{L}$ is decidable if there is a deterministic Turing machine that halts on every input $x$, returning “accept” if $x\in\mathcal{L}$ and “reject” otherwise. Since many types of objects can be encoded as strings, we often talk about sets of other objects being RE or coRE. For instance, the set
[…]
is the prototypical example of an undecidable RE set, although formalizing it as a language requires choosing some encoding of pairs $(M,x)$ as strings. The complement $\operatorname{coHALT}$ of pairs $(M,x)$ such that $M$ does not halt on input $x$ is the prototypical example of an undecidable coRE set. A function $f$ from strings on $\Sigma$ to strings on $\Sigma^{\prime}$ is computable if there is a deterministic Turing machine with input alphabet $\Sigma$ and output alphabet $\Sigma^{\prime}$ which takes input $x$ to $f(x)$. A language $\mathcal{L}$ of strings over $\Sigma$ is RE-hard (resp. coRE-hard) if there is a computable function $f$ from pairs of Turing machines $(M,x)$ to strings over $\Sigma$, such that $f(M,x)\in\mathcal{L}$ if and only if $M$ halts on input $x$ (sometimes this is called RE-hard under many-one reductions). The definition of RE and coRE does not depend on the choice of alphabet $\Sigma$, because strings over a finite set can be reencoded as strings over $\{0,1\}$. Strings over $\{0,1\}$ are called bit strings or binary strings. Natural numbers can be encoded as binary strings as well, so we can also talk about RE, coRE, RE-hard, and coRE-hard subsets of $\mathbb{N}$.
We also care about how many steps it takes a Turing machine to accept an input. The time function of a non-deterministic Turing machine $M$ is the function $T:\mathbb{N}\rightarrow\mathbb{N}$, where $T(n)$ is the smallest possible integer such that, given an input string $x$ of length $\leq n$, if $M$ halts on $x$, then some computation path of $M$ halts in $\leq T(n)$ steps. We use the convention that every Turing machine takes at least one step to halt, so $T(n)\geq 1$ for all $n$. We say that $\mathcal{L}$ can be recognized in time $T^{\prime}:\mathbb{N}\rightarrow\mathbb{N}$ if there is a Turing machine $M$ recognizing $\mathcal{L}$ with time function $T(n)\leq T^{\prime}(n)$ for all $n$. In particular, $\mathcal{L}$ can be recognized in polynomial time if it can be recognized in time $Cn^{k}$ for some $C,k\in\mathbb{N}$. Similarly, we say that $\mathcal{L}$ can be recognized in deterministic time $T^{\prime}:\mathbb{N}\rightarrow\mathbb{N}$ if there is a deterministic Turing machine $M$ recognizing $\mathcal{L}$ with time function $T(n)\leq T^{\prime}(n)$ for all $n$. A function $f$ is computable in time $T(n)$ if there is a deterministic Turing machine computing $f$ which halts in $\leq T(n)$ steps on all input strings of length $\leq n$.
The classes RE and coRE are the first level of a larger hierarchy of classes called the arithmetical hierarchy. Although we don’t use it in the rest of the paper, the class $\Pi_{2}^{0}$ mentioned in the introduction is in the second level of this hierarchy. A language $\mathcal{L}$ is in $\Pi_{2}^{0}$ if there is a Turing machine $M$ taking a pair of strings $x$ and $y$ as input, such that $x\in\mathcal{L}$ if and only if $M$ halts on input $(x,y)$ for all $y$. As discussed in the introduction, the Helton-McCullough Positivstellensatz states that an element $\alpha$ of an archimedean *-algebra $\mathcal{A}$ is positive if and only if $\alpha+\epsilon$ is a sum of squares for all $\epsilon>0$. When $\mathcal{A}$ is finitely presented, one can effectively search through all sums of squares. So deciding positivity for finitely presented archimedean algebras is in $\Pi_{2}^{0}$.
2.3. Group theory
In addition to the elementary definitions above, we use some more advanced concepts from group theory. Suppose $S$ is a finite set of generators for a group $G$, and let $\mathcal{L}$ be the set of words over $S\cup S^{-1}$ which are trivial in $G$. The word problem for $G$ is the problem of determining whether a word over $S\cup S^{-1}$ belongs to $\mathcal{L}$. Accordingly, we say that the word problem for $G$ with respect to $S$ is solvable in non-deterministic time $T:\mathbb{N}\rightarrow\mathbb{N}$ (resp. non-deterministic polynomial time) if $\mathcal{L}$ can be recognized in time $T$ (resp. polynomial time).
If $G$ has generating set $S$ (not necessarily finite), a normal form for $G$ with respect to $S$ is a function $\eta$ from words over $S\cup S^{-1}$ to words over $S\cup S^{-1}$, such that $\eta(x)=x$ in $G$ for all words $x$, and two words $x$ and $y$ are equal in $G$ if and only if $\eta(x)=\eta(y)$. Equivalently, a normal form can be specified as a function $\eta$ from $G$ to words over $S\cup S^{-1}$, such that $\eta(g)=g$ in $G$ for all $g\in G$.
We denote the free product of a family of groups $\{G_{i}\}_{i\in I}$ by $\Asterisk_{i\in I}G_{i}$, and the product by $\prod_{i\in I}G_{i}$. If $\{G_{i}\}_{i\in I}$ is a family of groups, $K$ is a group, and $\phi_{i}:K\rightarrow G_{i}$ are injective homomorphisms, then the amalgamated free product of the family over $K$ is
$_{K}\Asterisk_{i}G_{i}:=\langle\Asterisk_{i\in I}G_{i}:\phi_{i}(k)=\phi_{j}(k)\text{ for all }k\in K,i,j\in I\rangle.$
If $\phi_{i}(K)$ is in the centre of $G_{i}$ for all $i\in I$, then the central product of this family over $K$ is defined to be
$\sideset{{}_{K}}{{}_{i}}{\prod}G_{i}:=\langle\prod_{i}G_{i}:\phi_{i}(k)=\phi_{j}(k)\text{ for all }k\in K,i,j\in I\rangle.$
In both cases, if we identify $K$ with $\phi_{i}(K)$, then the image of $K$ inside $_{K}\Asterisk_{i}G_{i}$ and $\sideset{{}_{K}}{}{\prod}_{i}G_{i}$ is independent of what $i$ we pick. Note that given a generating set $S$ for $K$, we only need the relations $\phi_{i}(k)=\phi_{j}(k)$ for all $k\in S$ in the above presentations.
For each $i\in I$, let $A_{i}\subset G_{i}$ be a set of right coset representatives for $\phi_{i}(K)$, where we pick the identity element $1$ as the representative of $\phi_{i}(K)$ itself. The normal form theorem for amalgamated free products states that every element $g\in{}_{K}\Asterisk_{i}G_{i}$ can be written uniquely as $g=kg_{1}\cdots g_{m}$, where $k\in K$, and $g_{j}\in A_{i_{j}}\setminus\{1\}$ for some sequence $i_{1},\ldots,i_{m}\in I$ with $i_{j}\neq i_{j+1}$ for all $1\leq j<m$. While the normal form theorem for amalgamated free products does not give us a normal form of $g$ in the above sense, if $\eta$ is a normal form for $K$ with respect to generating set $S$, and $\eta_{i}$ is a normal form for $G_{i}$ with respect to generating set $S_{i}$, $i\in I$, then $g\mapsto\eta(k)\eta_{i_{1}}(g_{1})\cdots\eta_{i_{m}}(g_{m})$ is a normal form for $_{K}\Asterisk_{i}G_{i}$ with respect to generating set $S\sqcup\coprod_{i\in I}S_{i}$ (where $\coprod$ is the disjoint union). Similarly, if we pick an ordering for $I$, the normal form theorem for central products states that every element $g\in\sideset{{}_{K}}{}{\prod}_{i}G_{i}$ can be written uniquely as $kg_{1}\cdots g_{m}$, where $k\in K$, and $g_{i}\in A_{i_{j}}\setminus\{1\}$ for some increasing sequence $i_{1}<i_{2}<\ldots<i_{m}$ in $I$. Again, if $\eta$ is a normal form for $K$ and $\eta_{i}$ is a normal form for $G_{i}$, $i\in I$, then $g\mapsto\eta(k)\eta_{i_{1}}(g_{1})\cdots\eta_{i_{m}}(g_{m})$ is a normal form for $\sideset{{}_{K}}{}{\prod}_{i}G_{i}$. In both cases, the normal form theorem implies that the natural inclusions $G_{j}\hookrightarrow{}_{K}\Asterisk_{i}G_{i}$ and $G_{j}\hookrightarrow\sideset{{}_{K}}{}{\prod}_{i}G_{i}$ are injections for all $j$, so $G_{j}$ can be regarded as a subgroup of the amalgamated free product or central product. One application of amalgamated free products that we will use is the following lemma:
Lemma 2.1.
Let $G=\langle S:R\rangle$. For any $y\in S$, $G$ is a subgroup of
$H:=\langle G,s,t:y=st,s^{2}=t^{2}=1\rangle$
via the natural inclusion.
Proof.
Suppose $y$ has order $k\in\mathbb{N}\cup\{+\infty\}$. Then $H$ is the amalgamated free product of $G$ with the dihedral group
$D_{k}=\langle s,t:s^{2}=t^{2}=(st)^{k}=1\rangle$
of order $2k$ over the subgroups $\langle y\rangle\cong\mathbb{Z}_{k}\cong\langle s,t\rangle$. ∎
Another construction we’ll use are HNN extensions. Suppose that $G$ is a group, $H$ and $K$ are subgroups of $G$, and $\phi:H\to K$ is an isomorphism. Then the HNN extension of $G$ by $\phi$ is
$\langle G,t:tht^{-1}=\phi(h)\text{ for all }h\in H\rangle.$
It is well-known that the order of the generator $t$ is infinite, and that, similar to the amalgamated free product and central product, $G$ is a subgroup of the HNN extension via the natural inclusion. More generally, given a sequence of subgroups $H_{1},\ldots,H_{m}$ and $K_{1},\ldots,K_{m}$ of a group $G$, along with isomorphisms $\phi_{i}:H_{i}\rightarrow K_{i}$, $1\leq i\leq m$, $G$ will be subgroup of the iterated HNN extension
$\langle G,t_{1},\ldots,t_{m}:t_{i}ht_{i}^{-1}=\phi_{i}(h)\text{ for all }1\leq i\leq m\text{ and }h\in H_{i}\rangle.$
Britton’s lemma allows us to determine the trivial words in an iterated HNN extension.
Lemma 2.2 (Britton’s lemma, Theorem 11.81 of [ROT73]).
Let $H_{1},\ldots,H_{m}$ and $K_{1},\ldots,K_{m}$ be subgroups of a group $G$, and suppose $\phi_{i}:H_{i}\rightarrow K_{i}$ is an isomorphism for $i=1\ldots,m$. Let
$L=\langle G,t_{1},\ldots,t_{m}:t_{i}ht_{i}^{-1}=\phi_{i}(h)\text{ for all }1\leq i\leq m\text{ and }h\in H_{i}\rangle$
be the iterated HNN extension, and suppose
$w=g_{0}t_{i_{1}}^{e_{1}}g_{1}t^{e_{2}}_{i_{2}}\cdots t^{e_{n}}_{i_{n}}g_{n}\in L$
where $g_{0},\ldots,g_{n}\in G$, $e_{1},\ldots,e_{n}\in\{\pm 1\}$, and $1\leq i_{1},\ldots,i_{n}\leq m$. If $w=1$ in $L$, then at least one of the following holds:
  1. (a)
    $n=0$ and $g_{0}=1$ in $G$,
  2. (b)
    there is $1\leq j<n$ such that $e_{j}=1=-e_{j+1}$, $i_{j}=i_{j+1}$, and $g_{i}\in H_{i_{j}}$, or
  3. (c)
    there is $1\leq j<n$ such that $e_{j}=-1=-e_{j+1}$, $i_{j}=i_{j+1}$, and $g_{i}\in K_{i_{j}}$.
Let $G=\langle S:R\rangle$ be a finitely presented group. A function $f:\mathbb{N}\rightarrow\mathbb{N}$ is an isoperimetric function for $G$ if for all words $w=s_{1}\cdots s_{n}$ of length $n\geq 1$ over $S\cup S^{-1}$, if $w=1$ in $G$ then there are elements $u_{1},\ldots,u_{k}\in\mathcal{F}_{S}$ and $r_{1},\ldots,r_{k}\in R\cup R^{-1}$ with $k\leq f(n)$ and
$w=u_{1}r_{1}u_{1}^{-1}\cdot u_{2}r_{2}u_{2}^{-1}\cdots u_{k}r_{k}u_{k}^{-1}.$
Lemma 2.3 ([GER93]).
Let $G=\langle S:R\rangle$ be a finitely presented group with an isoperimetric function $f$, and let $l:=\max\{\lvert r\rvert:r\in R\}$ be the length of the longest relation in $R$. If a word $w\in\mathcal{F}_{S}$ of length $n$ is equal to $1$ in $G$, then there are $z_{1},\cdots,z_{k}\in\mathcal{F}_{S}$ and $r_{1},\cdots,r_{k}\in R\cup R^{-1}$ with $k\leq f(n)$ and $\lvert z_{i}\rvert\leq kl+l+n$ for all $1\leq i\leq k$ such that
$w=z_{1}r_{1}z_{1}^{-1}z_{2}r_{2}z_{2}^{-1}\cdots z_{k}r_{k}z_{k}^{-1}.$
in $\mathcal{F}_{S}$.
We use the following theorem of Birget, Ol’shanskii, Rips, and Sapir:
Theorem 2.4 ([BOR+02]).
Let $G$ be a group with finite generating set $S$. If the word problem for $G$ with respect to $S$ is solvable in non-deterministic time $T(n)$, where $T^{4}$ is superadditive, then $G$ can be embedded into a finitely presented group $H$ with isoperimetric function $n^{2}T(n^{2})^{4}$.
Here a function $T:\mathbb{N}\rightarrow\mathbb{N}$ is superadditive if $T(m+n)\geq T(m)+T(n)$ for all $m,n$. It follows immediately from this theorem that if the word problem for $G$ can be solved in non-deterministic polynomial time, then $G$ can be embedded in a finitely presented group $H$ with a polynomial isoperimetric function.
3. Approximate representations
We introduce some terminology for approximate states.
Definition 3.1.
Let $\mathcal{A}$ be a finitely generated *-algebra. Suppose $\epsilon\geq 0$ and $\mathcal{R}\subseteq\mathcal{A}$. An $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ is a state $\varphi$ on $\mathcal{A}$ such that $\varphi(r^{*}r)\leq\epsilon$ for all $r\in\mathcal{R}\cup\mathcal{R}^{*}$. If, in addition, $\varphi$ is tracial, then we say that $\varphi$ is a tracial $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$.
If $(\pi:\mathcal{A}\to\mathcal{B}(\mathcal{H}),\xi)$ is the GNS representation of a state $\varphi$ on $\mathcal{A}$, then $\lVert x\rVert_{\xi}:=\lVert x\cdot\xi\rVert$ is a semi-norm on $\mathcal{B}(\mathcal{H})$ such that $\lVert a\rVert_{\varphi}=\lVert\pi(a)\rVert_{\xi}$ for all $a\in\mathcal{A}$. If $\varphi$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ as in Definition 3.1, then $\lVert\pi(r)\rVert_{\xi}\leq\sqrt{\epsilon}$ for all $r\in\mathcal{R}$, so we can think of $\pi$ as a $\sqrt{\epsilon}$-approximate representation for $\mathbb{C}^{*}\langle\mathcal{A}:\mathcal{R}\rangle$ with respect to $\lVert\cdot\rVert_{\xi}$.
Definition 3.2.
Let $\mathcal{A}$ be a *-algebra with generating set $\mathcal{X}$. Let $\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ be a set of *-polynomials over $\mathcal{X}$. For any *-polynomial $f\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ that is trivial in $\mathcal{A}/\langle\mathcal{R}\rangle$, we say that $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$ is an $\mathcal{R}$-decomposition for $f$ in $\mathcal{A}$ if
  1. (1)
    $u_{i},v_{i}$ are *-monomials in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for all $1\leq i\leq n$,
  2. (2)
    $r_{i}\in\mathcal{R}\cup\mathcal{R}^{*}$ for all $1\leq i\leq n$,
  3. (3)
    $\lambda_{i}\in\mathbb{C}$ for all $1\leq i\leq n$, and
  4. (4)
    $f=\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$ in $\mathcal{A}$.
The size of an $\mathcal{R}$-decomposition $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$ is $\sum_{i=1}^{n}|\lambda_{i}|(1+\lVert r_{i}\rVert_{\mathcal{A}}\deg(v_{i}))$, where $\lVert\cdot\rVert_{\mathcal{A}}$ is the operator norm in $\mathcal{A}$.
Here $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ means the set of formal *-polynomials with indeterminates $\mathcal{X}$. There is a natural map from $\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{A}$ sending $x\in\mathcal{X}$ to its image in $\mathcal{A}$, so we can also regard elements of $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ as elements of $\mathcal{A}$, e.g. when taking the quotient $\mathcal{A}/\langle\mathcal{R}\rangle$. A *-monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a product $a_{1}a_{2}\cdots a_{k}$ where $k\geq 0$ and $a_{1},\ldots,a_{k}\in\mathcal{X}\cup\mathcal{X}^{*}$ (the integer $k$ is called the degree).
We let $\lVert\cdot\rVert_{1}$ and $\lVert\cdot\rVert_{1,1}$ denote the $1$-norm and first Sobolev $1$-seminorm on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ with respect to the basis of *-monomials. In other words, if $f=\sum_{i=1}^{n}\lambda_{i}u_{i}\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for distinct *-monomials $u_{1},\ldots,u_{n}$, then $\lVert f\rVert_{1}:=\sum_{i=1}^{n}|\lambda_{i}|$ and $\lVert f\rVert_{1,1}:=\sum_{i=1}^{n}|\lambda_{i}|\deg(u_{i})$. Note that $\lVert\cdot\rVert_{1}$ is submultiplicative, and that $\lVert f^{*}\rVert_{1}=\lVert f\rVert_{1}$, $\lVert f^{*}\rVert_{1,1}=\lVert f\rVert_{1,1}$ for all $f\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$. The following elementary lemma is useful for computing sizes of $\mathcal{R}$-decompositions.
Lemma 3.3.
Let $\mathcal{A}$ be a *-algebra with generating set $\mathcal{X}$, and let $\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ be a set of *-polynomials over $\mathcal{X}$. Then:
  1. (a)
    If $f=\sum_{i=1}^{n}u_{i}r_{i}v_{i}$ for *-polynomials $u_{1},\ldots,u_{n},v_{1}\ldots,v_{n}\in\mathbb{C}^{*}\langle X\rangle$ and $r_{1},\ldots,r_{n}\in\mathcal{R}\cup\mathcal{R}^{*}$, then $f$ has an $\mathcal{R}$-decomposition of size at most
    $\sum_{i=1}^{n}\lVert u_{i}\rVert_{1}\lVert v_{i}\rVert_{1}+\lVert u_{i}\rVert_{1}\lVert r_{i}\rVert_{\mathcal{A}}\lVert v_{i}\rVert_{1,1}.$
  2. (b)
    If fgf-g has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{1}$, and ghg-h has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{2}$, then fhf-h has an $\mathcal{R}$-decomposition of size $\leq\Lambda_{1}+\Lambda_{2}$.
  3. (c)
    If $g$ has an $\mathcal{R}$-decomposition of size $\leq\Lambda$, then $fg$ has an $\mathcal{R}$-decomposition of size $\leq\lVert f\rVert_{1}\Lambda$ for all $f\in\mathbb{C}^{*}\langle X\rangle$.
For instance, suppose we can demonstrate that $g=h$ in $\mathcal{A}/\langle\mathcal{R}\rangle$ by finding $f_{1},\ldots,f_{n}$, $u_{1},\ldots,u_{n}$, $v_{1},\ldots,v_{n}\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $r_{1},\ldots,r_{n}\in\mathcal{R}\cup\mathcal{R}^{*}$ such that $f_{i+1}-f_{i}=u_{i}r_{i}v_{i}$ for all $0\leq i\leq n$, where $f_{0}=g$ and $f_{n+1}=h$. Then we know that ghg-h has an $\mathcal{R}$-decomposition of size at most $\sum_{i=1}^{n}\lVert u_{i}\rVert_{1}\lVert v_{i}\rVert_{1}+\lVert u_{i}\rVert_{1}\lVert r_{i}\rVert_{\mathcal{A}}\lVert v_{i}\rVert_{1,1}$.
If $\mathcal{X}$ is a set of unitary generators for $\mathcal{A}$, then all *-monomials in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ are unitary in $\mathcal{A}$. As a result, an $\mathcal{R}$-decomposition of $f$ can be used to bound the state-dependent norm of $f$ with respect to any tracial $(\epsilon,\mathcal{R})$-state:
Proposition 3.4.
Suppose $\mathcal{A}$ is a *-algebra generated by a finite set of unitaries $\mathcal{X}$. Let $\mathcal{R}\subset\mathbb{C}^{*}\langle\mathcal{X}\rangle$ be a set of *-polynomials over $\mathcal{X}$, and let $\tau$ be a tracial $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$. If $f\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is trivial in $\mathcal{A}/\langle\mathcal{R}\rangle$ with an $\mathcal{R}$-decomposition $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$, then $\lVert f\rVert_{\tau}\leq(\sum_{i=1}^{n}|\lambda_{i}|)\sqrt{\epsilon}$.
Note that $\sum_{i=1}^{n}|\lambda_{i}|$ is less than or equal to the size of $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$.
Proof.
Since $\lVert\cdotp\rVert_{\tau}$ is unitarily invariant,
$\lVert f\rVert_{\tau}\leq\sum_{i=1}^{n}|\lambda_{i}|\lVert u_{i}r_{i}v_{i}\rVert_{\tau}=\sum_{i=1}^{n}|\lambda_{i}|\lVert r_{i}\rVert_{\tau}\leq(\sum_{i=1}^{n}|\lambda_{i}|)\sqrt{\epsilon}.$
The proof of Proposition 3.4 does not work if $\tau$ is not tracial. However, we can still bound $\lVert f\rVert_{\tau}$ using the size of an $\mathcal{R}$-decomposition if $\tau$ is the restriction of a state on $\mathcal{A}\otimes\mathcal{A}$ satisfying a synchronicity [^2] The term synchronous is borrowed from nonlocal games in quantum information. Roughly speaking, a state $\varphi$ is synchronous if the action of $x\in\mathcal{X}$ on $\varphi$ is the same in both tensor factors. condition:
Definition 3.5.
Suppose $\mathcal{A}$ is a *-algebra. A state $\varphi$ on $\mathcal{A}\otimes\mathcal{A}$ is $(\epsilon,\mathcal{X})$-synchronous for some $\epsilon\geq 0$ and $\mathcal{X}\subseteq\mathcal{A}$ if
(3.1) $\lVert x\otimes 1-1\otimes x\rVert_{\varphi}\leq\sqrt{\epsilon}$
for all $x\in\mathcal{X}$.
When working with $(\epsilon,\mathcal{X})$-synchronous states, it is convenient to let $\omega_{\mathcal{X}}$ denote the linear map $\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathbb{C}^{*}\langle\mathcal{X}\rangle$ sending $x_{1}x_{2}\cdots x_{n}\mapsto x_{n}x_{n-1}\cdots x_{1}$ for all $n\in\mathbb{N}$ and $x_{1},\ldots,x_{n}\in\mathcal{X}\cup\mathcal{X}^{*}$. This map satisfies $\omega_{\mathcal{X}}(a^{*})=\omega_{\mathcal{X}}(a)^{*}$ and $\omega_{\mathcal{X}}(ab)=\omega_{\mathcal{X}}(b)\omega_{\mathcal{X}}(a)$ for all $a,b\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and hence $\omega_{\mathcal{X}}$ is a *-isomomorphism $\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathbb{C}^{*}\langle\mathcal{X}\rangle^{\mathrm{op}}$, where $\mathbb{C}^{*}\langle\mathcal{X}\rangle^{\mathrm{op}}$ is the opposite algebra of $\mathbb{C}^{*}\langle\mathcal{X}\rangle$.
Lemma 3.6.
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$. If $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$, then
(3.2) $\lVert u\otimes 1-1\otimes\omega_{\mathcal{X}}(u)\rVert_{\varphi}\leq\lVert u\rVert_{1,1}\sqrt{\epsilon}$
for all $u\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$.
Proof.
For every $x\in\mathcal{X}$, since $x$ is unitary and $\lVert\cdotp\rVert_{\varphi}$ is left unitarily invariant,
$\lVert x^{*}\otimes 1-1\otimes x^{*}\rVert_{\varphi}=\lVert(x\otimes x)(x^{*}\otimes 1-1\otimes x^{*})\rVert_{\varphi}=\lVert x\otimes 1-1\otimes x\rVert_{\varphi}\leq\sqrt{\epsilon}.$
Thus Equation 3.1 holds for all $x\in\mathcal{X}\cup\mathcal{X}^{*}$. Suppose $u=x_{1}\cdots x_{n}$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ of degree $n\geq 1$. Let $x_{0}=x_{n+1}=1$. Then
$\lVert u\otimes 1-1\otimes\omega_{\mathcal{X}}(u)\rVert_{\varphi}$ $=\lVert\sum_{k=1}^{n}(x_{0}x_{1}\cdots x_{k-1}\otimes x_{n+1}x_{n}\cdots x_{k+1})(x_{k}\otimes 1-1\otimes x_{k})\rVert_{\varphi}$
$\leq\sum_{k=1}^{n}\lVert x_{k}\otimes 1-1\otimes x_{k}\rVert_{\varphi}\leq n\sqrt{\epsilon}.$
By the subadditivity of $\lVert\cdotp\rVert_{\varphi}$, Equation 3.2 holds for all $u\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$. ∎
The following lemmas show that some of the properties of tracial states hold approximately for $(\epsilon,\mathcal{X})$-synchronous states.
Lemma 3.7.
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$. If $u$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $a\in\mathcal{A}$, then $|\lVert u^{*}au\otimes 1\rVert_{\varphi}-\lVert a\otimes 1\rVert_{\varphi}|\leq\lVert a\rVert_{\mathcal{A}}\deg(u)\sqrt{\epsilon}$.
Proof.
Since $u$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, $u$ and $\omega_{\mathcal{X}}(u)$ are unitary in $\mathcal{A}$. By Lemma 3.6, $\lVert u\otimes 1-1\otimes\omega_{\mathcal{X}}(u)\rVert_{\varphi}\leq\deg(u)\sqrt{\epsilon}$. Thus
$\lVert u^{*}au\otimes 1\rVert_{\varphi}$ $\leq\lVert(u^{*}a\otimes 1)(u\otimes 1-1\otimes\omega_{\mathcal{X}}(u))\rVert_{\varphi}+\lVert u^{*}a\otimes\omega_{\mathcal{X}}(u)\rVert_{\varphi}$
$=\lVert(a\otimes 1)(u\otimes 1-1\otimes\omega_{\mathcal{X}}(u))\rVert_{\varphi}+\lVert a\otimes 1\rVert_{\varphi}$
$\leq\lVert a\rVert_{\mathcal{A}}\deg(u)\sqrt{\epsilon}+\lVert a\otimes 1\rVert_{\varphi}.$
Replacing $u$ with $u^{*}$ and $a$ with $u^{*}au$ in this calculation, we see that
$\lVert a\otimes 1\rVert_{\varphi}\leq\lVert a\rVert_{\mathcal{A}}\deg(u)\sqrt{\epsilon}+\lVert u^{*}au\otimes 1\rVert_{\varphi},$
and hence $|\lVert u^{*}au\otimes 1\rVert_{\varphi}-\lVert a\otimes 1\rVert_{\varphi}|\leq\lVert a\rVert_{\mathcal{A}}\deg(u)\sqrt{\epsilon}$. ∎
Lemma 3.8.
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$. If $u$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $a\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a hermitian square in $\mathcal{A}$, then
$\operatorname{Re}\big(\varphi(u^{*}aua\otimes 1)\big)\geq-\lVert u^{*}au\rVert_{1,1}\lVert a\otimes 1\rVert_{\varphi}\sqrt{\epsilon}.$
Proof.
Observe that
$\varphi(u^{*}aua\otimes 1)=\varphi\big(a\otimes\omega_{\mathcal{X}}(u^{*}au)\big)+\varphi\Big(\big(u^{*}au\otimes 1-1\otimes\omega_{\mathcal{X}}(u^{*}au)\big)\big(a\otimes 1\big)\Big).$
Since $a=b^{*}b$ in $\mathcal{A}$ for some $b\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$, $a\otimes\omega_{\mathcal{X}}(u^{*}au)=\big(b\otimes\omega_{\mathcal{X}}(u^{*}b^{*})\big)^{*}\big(b\otimes\omega_{\mathcal{X}}(u^{*}b^{*})\big)$ is a hermitian square in $\mathcal{A}$. Hence $\varphi\big(a\otimes\omega_{\mathcal{X}}(u^{*}au)\big)\geq 0$. By the Cauchy–Schwarz inequality and Lemma 3.6,
$\left|\varphi\Big(\big(u^{*}au\otimes 1-1\otimes\omega_{\mathcal{X}}(u^{*}au)\big)\big(a\otimes 1\big)\Big)\right|$ $\leq\lVert u^{*}au\otimes 1-1\otimes\omega_{\mathcal{X}}(u^{*}au)\rVert_{\varphi}\cdotp\lVert a\otimes 1\rVert_{\varphi}$
$\leq\lVert u^{*}au\rVert_{1,1}\lVert a\otimes 1\rVert_{\varphi}\sqrt{\epsilon}.$
Since $\operatorname{Re}(\alpha)\geq-\lvert\alpha\rvert$ for every complex number $\alpha$, the lemma follows. ∎
Lemma 3.9.
Let $\mathcal{A}$ be a *-algebra generated by a finite set of unitaries $\mathcal{X}$, and let $\iota:\mathcal{A}\to\mathcal{A}\otimes\mathcal{A}:a\mapsto a\otimes 1$ be the left inclusion. Suppose $\mathcal{R}\subseteq\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a set of *-polynomials over $\mathcal{X}$, and let $\varphi$ be an $(\epsilon,\mathcal{X})$-synchronous state on $\mathcal{A}\otimes\mathcal{A}$ such that $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$. If $f\in\mathcal{A}$ has an $\mathcal{R}$-decomposition of size $\Lambda$, then $\lVert f\rVert_{\tau}\leq\Lambda\sqrt{\epsilon}$.
Proof.
Let $\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}$ be an $\mathcal{R}$-decomposition for $f$ in $\mathcal{A}$ of size $\Lambda$. By Lemma 3.6, $\lVert v_{i}\otimes 1-1\otimes\omega_{\mathcal{X}}(v_{i})\rVert_{\varphi}\leq deg(v_{i})\sqrt{\epsilon}$ for every $1\leq i\leq n$. Since $\varphi\circ\iota$ is an $(\epsilon,\mathcal{R})$-state on $\mathcal{A}$ and $1\otimes\omega_{\mathcal{X}}(v_{i})$ is unitary in $\mathcal{A}\otimes\mathcal{A}$, it follows that $\lVert r_{i}\otimes\omega_{\mathcal{X}}(v_{i})\rVert_{\varphi}=\lVert r_{i}\rVert_{\varphi\circ\iota}\leq\sqrt{\epsilon}$ for all $1\leq i\leq n$. Then
$\lVert f\otimes 1\rVert_{\varphi}$ $=\lVert\sum_{i=1}^{n}\lambda_{i}u_{i}r_{i}v_{i}\otimes 1\rVert_{\varphi}\leq\sum_{i=1}^{n}\lvert\lambda_{i}\rvert\lVert r_{i}v_{i}\otimes 1\rVert_{\varphi}$
$\leq\sum_{i=1}^{n}\lvert\lambda_{i}\rvert\big(\lVert(r_{i}\otimes 1)(v_{i}\otimes 1-1\otimes\omega_{\mathcal{X}}(v_{i}))\rVert_{\varphi}+\lVert r_{i}\otimes\omega_{\mathcal{X}}(v_{i})\rVert_{\varphi}\big)$
$\leq\sum_{i=1}^{n}\lvert\lambda_{i}\rvert\big(\lVert r_{i}\rVert_{\mathcal{A}}\deg(v_{i})\sqrt{\epsilon}+\sqrt{\epsilon}\big)=\Lambda\sqrt{\epsilon},$
so $\lVert f\rVert_{\tau}=\lVert f\otimes 1\rVert_{\varphi}\leq\Lambda\sqrt{\epsilon}$. ∎
3.1. Dealing with infinite operator norm
The results presented so far in this section are useful for algebras $\mathcal{A}$ with $\lVert a\rVert_{\mathcal{A}}<\infty$ for all $a\in\mathcal{A}$. However, for the free algebras $\mathcal{A}=\mathbb{C}^{*}\langle x_{1},\cdots,x_{n}\rangle$ the operator norm $\lVert a\rVert_{\mathcal{A}}=\infty$ for all $a\in\mathcal{A}$ which are not scalar multiples of identity, making the size of an $\mathcal{R}$-decomposition infinite as well. To handle this situation, we show in this section that for any element $f$ in the free algebra $\mathcal{A}$, there is a set of relations $\mathcal{W}(f)$ such that if $\tau$ is a tracial $(\epsilon,\mathcal{W}(f))$-state on $\mathcal{A}$, then there is a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*n}$ which is close to $\tau$ on $f$.
Lemma 3.10.
For any $A\in\mathcal{B}(\mathcal{H})$, the operator $\widetilde{A}:=\text{sgn}\big(\frac{A+A^{*}}{2}\big)$ is a unitary involution in $C^{*}\langle A\rangle^{\prime\prime}$. If in addition, $\xi$ is a vector in $\mathcal{H}$ satisfying (i) $\lVert(A^{2}-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, (ii) $\lVert((A^{*})^{2}-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, (iii) $\lVert(AA^{*}-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, (iv) $\lVert(A^{*}A-\mathbbm{1}_{\mathcal{H}})\xi\rVert\leq\epsilon$, and (v) $\lVert(A-A^{*})\xi\rVert\leq\epsilon$ for some $\epsilon\geq 0$, then $\lVert(A-\widetilde{A})\xi\rVert\leq 2\epsilon$ and $\lVert(A^{*}-\widetilde{A})\xi\rVert\leq 2\epsilon$.
Here sgn denotes the function $\mathbb{R}\to\{-1,1\}$ sending elements in $0,+\infty)$ to $1$ and elements in $(-\infty,0)$ to 1-1.
Proof.
The spectrum of $\tfrac{A+A^{*}}{2}$ is in $\mathbb{R}$, so by Borel functional calculus, $\widetilde{A}:=\text{sgn}(\tfrac{A+A^{*}}{2})$ is a unitary involution and is in $C^{*}\langle A\rangle^{\prime\prime}$. Since sgn and $\lvert\cdotp\rvert$ are measurable functions on $\mathbb{R}$ satisfying $(1+\lvert a\rvert)(\text{sgn}(a)-a)=\text{sgn}(a)(1-a^{2})$ for all $a\in\mathbb{R}$, we obtain that
$\lVert(\widetilde{A}-\tfrac{A+A^{*}}{2})\xi\rVert$ $=\lVert(\text{sgn}(\tfrac{A+A^{*}}{2})-\tfrac{A+A^{*}}{2})\xi\rVert\leq\lVert(\operatorname{Id}_{\mathcal{H}}+\lvert\tfrac{A+A^{*}}{2}\rvert)(\text{sgn}(\tfrac{A+A^{*}}{2})-\tfrac{A+A^{*}}{2})\xi\rVert$
$=\lVert\text{sgn}(\tfrac{A+A^{*}}{2})(\operatorname{Id}_{\mathcal{H}}-(\tfrac{A+A^{*}}{2})^{2})\xi\rVert=\lVert(\operatorname{Id}_{\mathcal{H}}-(\tfrac{A+A^{*}}{2})^{2})\xi\rVert$
$=\lVert(\operatorname{Id}_{\mathcal{H}}-\tfrac{A^{2}+(A^{*})^{2}+AA^{*}+A^{*}A}{4})\xi\rVert\leq\epsilon,$
where the last inequality follows from the conditions (i)(iv)(i)-(iv). By condition $(v)$, $\lVert(A-\tfrac{A+A^{*}}{2})\xi\rVert=\lVert\tfrac{A-A^{*}}{2}\xi\rVert\leq\tfrac{\epsilon}{2}$, so $\lVert(\widetilde{A}-A)\xi\rVert\leq\lVert(\widetilde{A}-\tfrac{A+A^{*}}{2})\xi\rVert+\lVert(A-\tfrac{A+A^{*}}{2})\xi\rVert\leq\tfrac{3}{2}\epsilon\leq 2\epsilon$. Similarly we have $\lVert(\widetilde{A}-A^{*})\xi\rVert\leq 2\epsilon$. ∎
Lemma 3.11.
Let $\mathcal{X}$ be a finite set. There is a computable mapping $\mathcal{W}$ from $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ to finite subsets of $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ such that
  1. (a)
    $\mathcal{W}(f)\subseteq\ker q$ for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, where $q:\mathbb{C}^{*}\langle\mathcal{X}\rangle\to\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the quotient map, and
  2. (b)
    for any tracial state $\tau$ on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, there is a tracial state $\widehat{\tau}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ and $\epsilon\geq 0$, if $\tau$ is a tracial $(\epsilon,\mathcal{W}(f))$-state, then
    $\big|\lVert f\rVert_{\widehat{\tau}}-\lVert f\rVert_{\tau}\big|\leq 2\lVert f\rVert_{1,1}\sqrt{\epsilon}.$
Proof.
If $u=x_{1}\cdots x_{n}\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is a monomial of degree $n\geq 0$, then define
$\mathcal{W}(u):=$ $\left\{(x_{i}x_{i}^{*}-1)x_{i+1}\cdots x_{n}:1\leq i\leq n\right\}\cup\left\{(x_{i}^{*}x_{i}-1)x_{i+1}\cdots x_{n}:1\leq i\leq n\right\}$
$\cup$ $\left\{(x_{i}^{2}-1)x_{i+1}\cdots x_{n}:1\leq i\leq n\right\}\cup\left\{((x_{i}^{*})^{2}-1)x_{i+1}\cdots x_{n}:1\leq i\leq n\right\}$
$\cup$ $\left\{(x_{i}-x_{i}^{*})x_{i+1}\cdots x_{n}:1\leq i\leq n\right\}$
(so in particular, $\mathcal{W}(1)=\emptyset$). If $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ has support $u_{1},\ldots,u_{\ell}$, or in other words $f=\sum_{i=1}^{\ell}\lambda_{i}u_{i}\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ for distinct *-monomials $u_{1},\ldots,u_{\ell}$ and non-zero coefficients $\lambda_{1},\ldots,\lambda_{\ell}$, we define $\mathcal{W}(f):=\bigcup_{i=1}^{\ell}\mathcal{W}(u_{i})$. Clearly $\mathcal{W}$ is computable. If $u$ is a monomial, then $q(s^{*}s)=0$ for all $s\in\mathcal{W}(u)$, so the same is true for $\mathcal{W}(f)$.
Let $\big(\pi:\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{B}(\mathcal{H}),v\big)$ be the GNS representation of $\tau$. Note that $T\mapsto\langle v,Tv\rangle$ extends to a tracial state on $\pi(\mathbb{C}^{*}\langle\mathcal{X}\rangle)^{\prime\prime}$. For every $x\in\mathcal{X}$, let $\widehat{\pi}(x):=\text{sgn}\big(\tfrac{\pi(x)+\pi(x)^{*}}{2}\big)$. By Lemma 3.10, $\widehat{\pi}(x)$ is a unitary involution in $C^{*}\langle\pi(x)\rangle^{\prime\prime}$ for every $x\in\mathcal{X}$, so $\widehat{\pi}$ induces a *-representation of $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ on $\mathcal{H}$ and $\widehat{\tau}(a):=\langle v,\widehat{\pi}(a)v\rangle$ defines a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$.
Suppose $u=x_{1}\cdots x_{n}$ is a monomial. If $\tau(s^{*}s)\leq\epsilon$ for all $s\in\mathcal{W}(u)$, then $\lVert\pi(s)v\rVert\leq\sqrt{\epsilon}$ for all $s\in\mathcal{W}(u)$, and hence by Lemma 3.10,
$\lVert\big(\widehat{\pi}(x_{i})-\pi(x_{i})\big)\pi(x_{i+1}\cdots x_{n})v\rVert\leq 2\sqrt{\epsilon}$
for all $1\leq i\leq n$. It follows that
$\lVert\big(\widehat{\pi}(u)-\pi(u)\big)v\rVert$ $\leq\sum_{i=1}^{n}\lVert\widehat{\pi}(x_{1}\cdots x_{i-1})\big(\widehat{\pi}(x_{i})-\pi(x_{i})\big)\pi(x_{i+1}\cdots x_{n})v\rVert$
$=\sum_{i=1}^{n}\lVert\big(\widehat{\pi}(x_{i})-\pi(x_{i})\big)\pi(x_{i+1}\cdots x_{n})\psi\rVert\leq 2n\sqrt{\epsilon}.$
For a general element $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, write $f=\sum_{i=1}^{\ell}\lambda_{i}u_{i}$ as above. If $\tau(s^{*}s)\leq\epsilon$ for all $s\in\mathcal{W}(f)$, then
[…]
$\leq\sum_{i=1}^{\ell}\lvert\lambda_{i}\rvert\cdotp\lVert\big(\widehat{\pi}(u_{i})-\pi(u_{i})\big)v\rVert\leq\sum_{i=1}^{\ell}2\lvert\lambda_{i}\rvert\deg(u_{i})\sqrt{\epsilon}=2\lVert f\rVert_{1,1}\sqrt{\epsilon}.$
We conclude that $\mathcal{W}$ is a computable mapping satisfying (a) and (b). ∎
Note that the proof of Lemma 3.11 only requires $\tau(s^{*}s)\leq\epsilon$ for all $s\in\mathcal{W}(f)$. We use the stronger hypothesis that $\tau$ is an $(\epsilon,\mathcal{W}(f))$-state for convenience.
Lemma 3.12.
Let $\mathcal{X}$ be a finite set, and let $\mathcal{W}$ be the map from Lemma 3.11. Suppose $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ for some $\epsilon\geq 0$, and $\varphi(s^{*}s\otimes 1)\leq\epsilon,\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$. Then there is a $(25\epsilon,\mathcal{X})$-synchronous state $\widehat{\varphi}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that for all $f\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, if $\varphi(s^{*}s\otimes 1)\leq\epsilon$ for all $s\in\mathcal{W}(f)$, then
$\big|\lVert f\otimes 1\rVert_{\widehat{\varphi}}-\lVert f\otimes 1\rVert_{\varphi}\big|\leq 2\lVert f\rVert_{1,1}\sqrt{\epsilon}.$
Proof.
Let $(\pi:\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle\rightarrow\mathcal{B}(\mathcal{H}),\psi)$ be the GNS representation of $\varphi$. For every $x\in\mathcal{X}$, let $\widehat{\pi}(x\otimes 1):=\text{sgn}\big(\tfrac{\pi(x\otimes 1)+\pi(x\otimes 1)^{*}}{2}\big)$ and $\widehat{\pi}(1\otimes x):=\text{sgn}\big(\tfrac{\pi(1\otimes x)+\pi(1\otimes x)^{*}}{2}\big)$. Then by Lemma 3.10, for every $x\in\mathcal{X}$, $\widehat{\pi}(x\otimes 1)$ and $\widehat{\pi}(1\otimes x)$ are unitary involutions in $C^{*}\langle\pi(x\otimes 1)\rangle^{\prime\prime}$ and $C^{*}\langle\pi(1\otimes x)\rangle^{\prime\prime}$ respectively. So for every $x,x^{\prime}\in\mathcal{X}$, $\pi(x\otimes 1),\pi(1\otimes x^{\prime})]=0$ implies $\widehat{\pi}(x\otimes 1),\widehat{\pi}(1\otimes x^{\prime})]=0$, and hence $\widehat{\pi}$ induces a *-representation of $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ on $\mathcal{H}$ and $\widehat{\varphi}(a):=\langle\psi,\widehat{\pi}(a)\psi\rangle$ defines a state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By the assumption, $\varphi(s^{*}s\otimes 1)\leq\epsilon$ and $\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$. This implies $\lVert\pi(s\otimes 1)\psi\rVert\leq\sqrt{\epsilon}$ and $\lVert\pi(1\otimes s)\psi\rVert\leq\sqrt{\epsilon}$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$, and hence by Lemma 3.10, $\lVert\big(\widehat{\pi}(x\otimes 1)-\pi(x\otimes 1)\big)\psi\rVert\leq 2\sqrt{\epsilon}$ and $\lVert\big(\widehat{\pi}(1\otimes x)-\pi(1\otimes x)\big)\psi\rVert\leq 2\sqrt{\epsilon}$ for all $x\in\mathcal{X}$. It follows that
$\lVert\big(\widehat{\pi}(x\otimes 1)-\widehat{\pi}(1\otimes x)\big)\psi\rVert\leq$ $\lVert\big(\widehat{\pi}(x\otimes 1)-\pi(x\otimes 1)\big)\psi\rVert+\lVert\big(\pi(x\otimes 1)-\pi(1\otimes x)\big)\psi\rVert$
$+\lVert\big(\widehat{\pi}(1\otimes x)-\pi(1\otimes x)\big)\psi\rVert\leq 2\sqrt{\epsilon}+\sqrt{\epsilon}+2\sqrt{\epsilon}=5\sqrt{\epsilon}.$
This means $\widehat{\varphi}$ is a $(25\epsilon,\mathcal{X})$-synchronous state $\widehat{\varphi}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. The rest of the lemma follows from a similar argument to the proof of Lemma 3.11. ∎
4. Construction of a suitable family of algebras
We begin the proofs of Theorems 1.1 and 1.3, with the case $\mathcal{A}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$. To explain the idea of the proof, suppose that $\mathcal{L}\subseteq\mathbb{N}$ is an RE set, and that we can find an integer $N$, a self-adjoint element $w\in\mathbb{Q}\mathbb{Z}_{2}^{*N}$, and a computable family of finitely presented algebras $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*N}:\mathcal{R}_{m}\rangle$ such that $\mathcal{A}_{\mathcal{L}}(m)$ has a tracial state $\tau$ with $\tau(w^{2})>0$ if and only if $m\in\mathcal{L}$. If $\alpha(m):=\sum_{r\in\mathcal{R}_{m}}r^{*}r-w^{2}$, then a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ with $\tau(w^{2})>0$ pulls back to a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ with $\tau(\alpha(m))<0$. Unfortunately this idea runs into a stability problem: it is not necessarily the case that every tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ with $\tau(\alpha(m))<0$ comes from a tracial state on $\mathcal{A}_{\mathcal{L}}(m)$. Indeed, it might be possible to find a tracial state $\tau$ for which $\tau\left(\sum_{r\in\mathcal{R}_{m}}r^{*}r\right)$ is small, but $\tau(w^{2})$ is large.
The main technical idea of this paper is that we can overcome the stability problem with a carefully chosen algebra $\mathcal{A}_{\mathcal{L}}(m)$. Specifically, we start with the algebra generated by a family of positive contractions $\{\widetilde{P}_{0},\widetilde{P}_{1},\ldots\}$ and a family of unitary involutions $\{\widetilde{X}_{0},\widetilde{X}_{1},\ldots\}$ satisfying the relations
(4.1) $\widetilde{P}_{n+1}=\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}$
for all $n$ such that $M_{\mathcal{L}}$ does not halt on input $m$ after $n$ steps, where $M_{\mathcal{L}}$ is a Turing machine recognizing $\mathcal{L}$. This algebra is not finitely presented, so the purpose of this section is to show that this algebra can be embedded in a finitely presented algebra $\mathcal{A}_{\mathcal{L}}(m)$, in such a way that the size of relation (4.1) is polynomial in $n$. Relation (4.1) implies that $\tau(\widetilde{P}_{n})=\tfrac{1}{2}\tau(\widetilde{P}_{n+1})$ for any tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$, and in the next section we use this fact to overcome the stability problem and finish the proofs of Theorems 1.1 and 1.3 (for the case $\mathcal{A}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$) using the strategy outlined in the previous paragraph.
Theorem 4.1.
Let $\mathcal{L}\subset\mathbb{N}$ be an RE set, and let $M_{\mathcal{L}}$ be a Turing machine that recognizes $\mathcal{L}$. Then there is a finite set $\mathcal{X}$, two sequences $\{\widetilde{P}_{0},\widetilde{P}_{1},$...\} and $\{\widetilde{X}_{0},\widetilde{X}_{1},$...\} in $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, and a computable mapping $m\mapsto\mathcal{R}_{m}$ from $\mathbb{N}$ to finite subsets of $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$, such that
  1. (a)
    for all $n\geq 0$, $\widetilde{P}_{n}$ is a hermitian square in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ with $\lVert\widetilde{P}_{n}\rVert_{1}=1$ and $\lVert\widetilde{P}_{n}\rVert_{1,1}=3(n+1)/2$, and $\widetilde{X}_{n}$ is a monomial in $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ of degree $4n+1$ such that $\widetilde{X}_{n}^{2}=1$ in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$,
  2. (b)
    there are constants $C>0,k\in\mathbb{N}$ such that for every $m\in\mathbb{N}$ and $n\geq 0$, if $M_{\mathcal{L}}$ hasn’t halted after $n$ steps on input $m$, then $\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}$ is trivial in $\mathcal{A}_{\mathcal{L}}(m):=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$ and has an $\mathcal{R}_{m}$-decomposition in $\mathbb{C}\mathbb{Z}_{2}^{\mathcal{X}}$ of size $\leq C\big((n+1)m\big)^{k}$,
  3. (c)
    if $m\in\mathcal{L}$, then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$, and
  4. (d)
    there is a positive integer $D$ such that $\lVert r\rVert_{1,1}\leq Dm$ for all $r\in\mathcal{R}_{m},m\in\mathbb{N}$.
In part (a), $\lVert\cdot\rVert_{1}$ and $\lVert\cdot\rVert_{1,1}$ denote the 1-norm and Sobolev seminorm on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ defined in the previous section. Since monomials in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ are unitary in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, $\lVert a\rVert_{\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}}\leq\lVert a\rVert_{1}$ for all $a\in\mathbb{C}^{*}\langle\mathcal{X}\rangle$. So part (a) implies that $\widetilde{P}_{n}$ is a contraction in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, i.e. $\lVert\widetilde{P}_{n}\rVert_{\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}}\leq 1$, and $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for all $n\geq 0$.
For the rest of this section, we fix an RE set $\mathcal{L}\subset\mathbb{N}$, and a deterministic Turing machine $M_{\mathcal{L}}$ that recognizes $\mathcal{L}$. Define $h:\mathbb{Z}\rightarrow\mathbb{N}\cup\{\infty\}$ by
[…]
so $h|_{\mathbb{N}}$ is the running time of $M_{\mathcal{L}}$. Recall that in our convention, all Turing machines take at least one step, so $h(m)\geq 1$ for all $m\in\mathbb{Z}$. To prove Theorem 4.1, we start in Section 4.1 by encoding the Turing machine $M_{\mathcal{L}}$ in certain group relations. We finish the proof by constructing $\mathcal{A}_{\mathcal{L}}(m)$ in Section 4.2.
4.1. Group stage of construction
In the first stage of constructing $\mathcal{A}_{\mathcal{L}}(m)$, we build up a group $G_{\mathcal{L}}$ out of amalgamations, central products, and HNN extensions. We show that $G_{\mathcal{L}}$ has a finite generating set, and that the word problem with respect to this generating set is decidable in polynomial time. This allows us to embed $G_{\mathcal{L}}$ in a finitely presented group $H_{S}$, which we use as the basis for defining the algebras $\mathcal{A}_{\mathcal{L}}(m)$.
For the construction of $G_{\mathcal{L}}$, let $A,B]:=ABA^{-1}B^{-1}$ denote the group commutator, and let $\mathbb{Z}_{n}$ be the additive cyclic group of order $n\in\mathbb{N}\cup\{+\infty\}$ (so in particular, $\mathbb{Z}_{\infty}=\mathbb{Z}$). To order the elements of $\mathbb{Z}_{n}$, we represent elements of $\mathbb{Z}_{n}$ by integers $\lfloor\tfrac{n+2}{2}\rfloor<j<\lfloor\tfrac{n+1}{2}\rfloor$. For instance, for $\mathbb{Z}_{4}$ we use representatives 2,1,0,1-2,-1,0,1. This choice of representatives has the property that every congruence class is represented by an element that is minimal in absolute value, i.e. if $k^{\prime}$ is the representative of $[k]$, then $|k^{\prime}|\leq|k|$. Also, note that this set of representatives contains 1-1 for every $n\geq 2$. When $n=+\infty$, elements of $\mathbb{Z}_{n}=\mathbb{Z}$ represent themselves.
For each $m\in\mathbb{Z}$ and $i\in\mathbb{Z}_{h(m)+1}$, let $K_{\mathcal{L}}(m,i)$ be the group generated by indeterminates $J,x_{mi},z_{mi}$, subject to the relations:
  • $J^{2}=x_{mi}^{2}=z_{mi}^{2}=1$,
  • $x_{mi},J]=[z_{mi},J]=1$, and
  • $x_{mi},z_{mi}]=J^{a}$, where $a=1$ if $i{+}1\neq 0$ in $\mathbb{Z}_{h(m)+1}$, and $a=0$ otherwise.
If $i+1\neq 0$ in $\mathbb{Z}_{h(m)+1}$, then the map $x_{mi}\mapsto a$, $z_{mi}\mapsto b$, $J\mapsto abab$ gives an isomorphism from $K_{S}(m,i)$ to the rank 4 dihedral group
$D_{4}=\langle a,b:a^{2}=b^{2}=(ab)^{4}=1\rangle,$
while if $i+1=0$ in $\mathbb{Z}_{h(m)+1}$ then all generators commute. Hence
$K_{\mathcal{L}}(m,i)=\begin{cases}D_{4}&i+1\neq 0\text{ in }\mathbb{Z}_{h(m)+1}\\ \mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}&i+1=0\text{ in }\mathbb{Z}_{h(m)+1}\end{cases}.$
In particular, $J$ is central and $\langle J\rangle=\mathbb{Z}_{2}$ in both cases. For $m\in\mathbb{Z}$, let
$K_{\mathcal{L}}(m):=\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{L}}(m,i)$
be the central product of the groups $K_{\mathcal{L}}(m,i)$, $i\in\mathbb{Z}_{h(m)+1}$, over the central subgroup $\langle J\rangle=\mathbb{Z}_{2}$, and let
$K_{\mathcal{L}}:={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{L}}(m)$
be the amalgamated free product of the groups $K_{\mathcal{L}}(m)$, $m\in\mathbb{Z}$, over $\langle J\rangle=\mathbb{Z}_{2}$. In other words, $K_{\mathcal{L}}$ is generated by indeterminates
$\mathcal{X}_{K}:=\{J,x_{mi},z_{mi},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}\},$
subject to relations
  1. (K0)
    $J^{2}=x_{mi}^{2}=z_{mi}^{2}=1$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$,
  2. (K1)
    $x_{mi},J]=[z_{mi},J]=1$ for all $m\in\mathbb{Z}$, $i\in\mathbb{Z}_{h(m)+1}$,
  3. (K2)
    for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$,
    $x_{mi},z_{mi}]=\begin{cases}J&i+1\neq 0\text{ in }\mathbb{Z}_{h(m)+1}\\ 1&i+1=0\text{ in }\mathbb{Z}_{h(m)+1}\end{cases},\text{ and }$
  4. (K3)
    for all $m\in\mathbb{Z}$ and $i\neq j\in\mathbb{Z}_{h(m)+1}$,
    $x_{mi},z_{mj}]=[x_{mi},x_{mj}]=[z_{mi},z_{mj}]=1.$
Note from relations (K1) that $J$ is central in $K_{\mathcal{L}}$. For each $m\in\mathbb{Z}$ and $i\in\mathbb{Z}_{h(m)+1}$, the elements $\{1,x_{mi},z_{mi},x_{mi}z_{mi}\}$ form a set of right coset representatives for $\langle J\rangle$. By the normal form theorem for central products, every element $w\in K_{\mathcal{L}}(m)$, $m\in\mathbb{Z}$ can be written uniquely as
(4.2) $w=J^{c}x_{m,i_{1}}^{a_{1}}z_{m,i_{1}}^{b_{1}}x_{m,i_{2}}^{a_{2}}z_{m,i_{2}}^{b_{2}}\cdots x_{m,i_{k}}^{a_{k}}z_{m,i_{k}}^{b_{k}},$
where $k\geq 0$, $i_{1}<i_{2}<\ldots<i_{k}$ are representatives of elements from $\mathbb{Z}_{h(m)+1}$ as described above, $c\in\mathbb{Z}_{2}$, and $(a_{j},b_{j})\in\{(0,1),(1,0),(1,1)\}\subset\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for all $1\leq j\leq k$. The function $\eta_{m}(w)=J^{c}x_{m,i_{1}}^{a_{1}}z_{m,i_{1}}^{b_{1}}\cdots x_{m,i_{k}}^{a_{k}}z_{m,i_{k}}^{b_{k}}$ is a normal form for $K_{\mathcal{L}}(m)$. Note that we take the indices $i_{1},\ldots,i_{k}$ to be from the chosen set of representatives for $\mathbb{Z}_{h(m)+1}$, rather than elements of $\mathbb{Z}_{h(m)+1}$, so $\eta(w)$ is a word over
$\widetilde{\mathcal{X}}_{K}:=\left\{J,x_{mi},z_{mi},m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\right\}$
rather than $\mathcal{X}_{K}$. We use this convention so that the indices are ordered, and so we can encode words over the generators as binary strings. For convenience in what follows, we assume that words over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$ are encoded in a fairly typical way as a sequence of substrings representing symbols and indices, where symbols are encoded with a fixed number of bits, and indices are written in binary with a sign bit. For instance, we might encode $x^{-1}_{m,i}$ by the bit sequence for $x^{-1}$ followed by the binary representations of $m$ and $i$ separated by delimiters, and $x^{-3}_{m,i}z_{m^{\prime},i^{\prime}}$ by this same sequence repeated three times (we don’t represent exponents in binary) followed by the encoding of $z_{m^{\prime},i^{\prime}}$. The bit-length (as opposed to the usual length) of a word over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$ is the length of this binary encoding. The bit-length of a word $w$ depends on the indices $m$ and $i$ of generators $x_{mi}^{\pm 1}$ and $z_{mi}^{\pm 1}$ appearing in $w$. In what follows, we need to pay special attention to how large $i$ gets, so we define the maximum index of a word $w$ over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$ to be
[…]
Returning to normal forms, the subset
$\mathcal{C}(m):=\big\{x_{m,i_{1}}^{a_{1}}z_{m,i_{1}}^{b_{1}}\cdots x_{m,i_{k}}^{a_{k}}z_{m,i_{k}}^{b_{k}}:\$$k\geq 0,-\lfloor\tfrac{h(m)+3}{2}\rfloor<i_{1}<\ldots<i_{k}<\lfloor\tfrac{h(m)+2}{2}\rfloor,$
$(a_{j},b_{j})\in\{(0,1),(1,0),(1,1)\},1\leq j\leq k\big\}$
is a set of right coset representatives for $\langle J\rangle$ in $K_{\mathcal{L}}(m)$. By the normal form theorem for amalgamated free products, every element $w\in K_{\mathcal{L}}$ can be written uniquely as
(4.3) $w=J^{c}w_{1}\cdots w_{\ell}$
where $\ell\geq 0$, $c\in\mathbb{Z}_{2}$, and $w_{j}\in\mathcal{C}(m_{j})\setminus\{1\}$ for some integers $m_{1},\ldots,m_{\ell}$ such that $m_{j}\neq m_{j+1}$ for all $1\leq j\leq\ell-1$. The function $\eta(w)=J^{c}w_{1}\cdots w_{\ell}$ is a normal form for $K_{\mathcal{L}}$. This normal form can be computed efficiently:
Proposition 4.2.
Let $\eta$ be the normal form for $K_{S}$ described above. If $w$ is given as a word over $\widetilde{X}_{K}\cup\widetilde{X}_{K}^{-1}$, then $\eta(w)$ is computable in polynomial time in the bit-length of $w$. If we write $\eta(w)=J^{c}w_{1}\cdots w_{\ell}$ as in Eq. 4.3, then the length, bit-length, and maximum index of $w_{1}\cdots w_{\ell}$ are at most the length, bit-length, and maximum index of $w$ respectively.
Proof.
To compute $\eta(w)$, suppose first that $w$ is a word over $\widetilde{\mathcal{X}}_{K}(m)\cup\widetilde{\mathcal{X}}_{K}(m)^{-1}$, where
$\widetilde{\mathcal{X}}_{K}(m):=\left\{J,x_{mi},z_{mi},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\right\}.$
Note that if $\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor$, then $i+1=0$ in $\mathbb{Z}_{h(m)+1}$ if and only if i=1i=-1. Using relations (K1)-(K3), we can reorder $w$ so that it can be written as in Eq. 4.2 for some $c,a_{1},b_{1},\ldots,a_{k},b_{k}\in\mathbb{Z}$ (rather than $\mathbb{Z}_{2}$). Using relations (K0), we can then reduce $c,a_{1},b_{1},\ldots,a_{k},b_{k}$ mod $2$ to put $w$ in normal form $\eta_{m}(w)=J^{c}\overline{w}$, where $\overline{w}\in\mathcal{C}(m)$. Since a list of integers can be sorted in polynomial time, this process takes polynomial time in the bit-length of $w$. Since every generator in $\overline{w}$ occurs in $w$, and no generator appears in $\overline{w}$ more than once, the length, bit-length, and maximum index of $\overline{w}$ are at most the length, bit-length, and maximum index of $w$ respectively.
If $w$ is a word over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$, then we can write $w=J^{c}w_{1}\cdots w_{\ell}$, where $c\in\mathbb{Z}_{2}$, $w_{j}$ is a word over $\widetilde{\mathcal{X}}_{K}(m_{j})\cup\widetilde{\mathcal{X}}_{K}(m_{j})^{-1}$ for $1\leq i\leq\ell$, and $m_{1},\ldots,m_{\ell}$ is some sequence of integers such that $m_{j}\neq m_{j+1}$ for $1\leq j\leq\ell-1$. Using the procedure in the last paragraph, we can put $w_{j}$ in normal form $J^{c_{j}}\overline{w}_{j}$ for $K_{\mathcal{L}}(m_{j})$, where $\overline{w}_{j}\in\mathcal{C}(m_{j})$ and $c_{j}\in\mathbb{Z}_{2}$. The length of $\overline{w}_{j}$ is at most the length of $w_{j}$, so the length of $\overline{w}_{1}\cdots\overline{w}_{\ell}$ is at most the length of $w_{1}\cdots w_{\ell}$, and the same is true of bit-length and maximum index. Let $c^{\prime}=c+\sum_{j}c_{j}$. If $\overline{w}_{j}\neq 1$ for all $1\leq j\leq\ell$, then $\eta(w)=J^{c^{\prime}}\overline{w}_{1}\cdots\overline{w}_{\ell}$. Otherwise, if $\overline{w}_{j}=1$ for some $1\leq j\leq\ell$, then we can repeat this procedure with $w^{\prime}=J^{c^{\prime}}\overline{w}_{1}\cdots\overline{w}_{j-1}\overline{w}_{j+1}\cdots\overline{w}_{\ell}$. Since $\ell$ decreases in each iteration of this algorithm, the overall running time will be polynomial in the bit-length of $w$. Also, since the length of $w_{1}\cdots w_{\ell}$ does not increase in each iteration, the length of $\overline{w}_{1}\cdots\overline{w}_{\ell}$ in the final iteration is at most the length of the initial input $w$, and the same is true of bit-length and maximum index. ∎
Remark 4.3.
Since, for instance, $\eta(z_{m0}x_{m0})=Jx_{m0}z_{m0}$, the length and bit-length of $\eta(w)$ is not necessarily less than the length and bit-length of $w$. Proposition 4.2 does imply that the length of $\eta(w)$ is at most one more than the length of $w$. Since the generator $J$ does not contribute to the maximum index, the maximum index of $\eta(w)$ is at most the maximum index of $w$.
Remark 4.4.
The above algorithm for computing $\eta$ assumes that the generators $x_{mi}$ and $z_{mi}$ in the input are given by specifying the representative of the index $i\in\mathbb{Z}_{h(m)+1}$. The algorithm does not work as stated without this assumption. For instance, on $w=x_{m,-3}\cdot x_{m,7}$ the algorithm will return $w$ back again, but if $h(m)=9$ then $x_{m,7}=x_{m,-3}$ in $K_{\mathcal{L}}$, and $\eta(w)=1$. If $i\in\mathbb{Z}$, then the representative for $i$ in $\mathbb{Z}_{h(m)+1}$ can be computed by running the Turing machine $M_{\mathcal{L}}$ on input $m$ for 2|i|12|i|-1 steps. If $M_{\mathcal{L}}$ halts by this point then we know $h(m)$ and can replace $i$ by its representative if necessary, while if it does not halt then h(m)>2|i|1h(m)>2|i|-1 and hence $\lfloor\frac{h(m)+3}{2}\rfloor<i<\lfloor\frac{h(m)+2}{2}\rfloor$. The running time of this procedure is polynomial in the bit-length of $m$ and the value of $i$, but not polynomial in the bit-length of $i$. We will use this procedure for computing the representative of $i$ later.
For the final step of constructing $G_{\mathcal{L}}$, we take iterated HNN extensions over several subgroups of $K_{\mathcal{L}}$:
$K_{\mathcal{L},x}$ $:=\langle J,x_{mi},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}\rangle,$
$K_{\mathcal{L},z}$ $:=\langle J,z_{mi},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}\rangle,\text{ and}$
$K_{\mathcal{L},0}$ $:=\langle J,x_{m,0},z_{m,0},m\in\mathbb{Z}\rangle.$
Lemma 4.5.
  1. (a)
    An element $w\in K_{\mathcal{L}}$ is in $K_{\mathcal{L},x}$ if and only if its normal form $\eta(w)$ is a word over $\{J,x_{mi},\ m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$. In addition, there is an automorphism $\phi_{S}:K_{\mathcal{L},x}\rightarrow K_{\mathcal{L},x}$ sending $J\mapsto J$ and $x_{mi}\mapsto x_{m,i+1}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$.
  2. (b)
    An element $w\in K_{\mathcal{L}}$ is in $K_{\mathcal{L},z}$ if and only if its normal form $\eta(w)$ is a word over $\{J,z_{mi},\ m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$. In addition, there is an automorphism $\phi_{T}:K_{\mathcal{L},z}\rightarrow K_{\mathcal{L},z}$ sending $J\mapsto J$ and   $z_{mi}\mapsto z_{m,i+1}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$.
  3. (c)
    An element $w\in K_{\mathcal{L}}$ is in $K_{\mathcal{L},0}$ if and only if its normal form $\eta(w)$ is a word over $\{J,x_{m,0},z_{m,0},m\in\mathbb{Z}\}$. In addition, there is an automorphism $\phi_{W}:K_{\mathcal{L},0}\rightarrow K_{\mathcal{L},0}$ sending $J\mapsto J$, $x_{m,0}\mapsto x_{m+1,0}$, and $z_{m,0}\mapsto z_{m+1,0}$ for all $m\in\mathbb{Z}$.
Proof.
For part (a), observe that the subgroup $K_{\mathcal{L},x}(m,i):=\langle J,x_{mi}\rangle\subseteq K_{\mathcal{L}}(m,i)$ is isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for all $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$. For $m\in\mathbb{Z}$, let $K_{\mathcal{L},x}(m):=\langle J,x_{mi}:i\in\mathbb{Z}_{h(m)+1}\rangle\subseteq K_{\mathcal{L}}(m)$, and consider the natural surjective homomorphism
$\Phi_{m}:\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{L},x}(m,i)\rightarrow K_{\mathcal{L},x}(m).$
By the normal form theorem for central products, every element $w$ in the domain of $\Phi_{m}$ can be written uniquely as $w=J^{c}x_{m,i_{1}}x_{m,i_{2}}\cdots x_{m,i_{k}}$, where $k\geq 0$, $\lfloor\tfrac{h(m)+3}{2}\rfloor<i_{1}<\cdots<i_{k}<\lfloor\tfrac{h(m)+2}{2}\rfloor$, and $c\in\mathbb{Z}_{2}$. Comparing with the normal form $\eta_{m}$ for $K_{\mathcal{L}}(m)$, we see that $\Phi_{m}$ sends normal forms to normal forms. It follows that $\Phi_{m}$ is an isomorphism, and that if $w\in K_{\mathcal{L},x}(m)$, then the normal form $\eta_{m}(w)$ in $K_{\mathcal{L}}(m)$ is a word containing only $J$ and $x_{mi}$’s.
A similar argument using the normal form theorem for amalgamated free products shows that
$K_{\mathcal{L},x}={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{L},x}(m),$
and the normal form $\eta(w)$ of an element $w\in K_{\mathcal{L},x}$ is a word over $\{J,x_{mi}:m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\}$.
For the automorphism, observe that
$\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{L},x}(m,i)\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{\oplus h(m)+1},$
where the first $\mathbb{Z}_{2}$ is generated by $J$, and $\mathbb{Z}_{2}^{\oplus h(m)+1}$ is generated by $\{x_{mi}:i\in\mathbb{Z}_{h(m)+1}\}$. Hence for every $m\in\mathbb{Z}$ there is an automorphism $\phi_{S}^{(m)}:K_{\mathcal{L},x}(m)\rightarrow K_{\mathcal{L},x}(m)$ sending $J\mapsto J$ and $x_{mi}\mapsto x_{m,i+1}$ for all $i\in\mathbb{Z}_{h(m)+1}$, and these extend to the desired automorphism $\phi_{S}:K_{\mathcal{L},x}\rightarrow K_{\mathcal{L},x}$.
The argument for part (b) is identical. For part (c), observe that $K_{\mathcal{L}}(m,0)\cong D_{4}$ for all $m\in\mathbb{Z}$ (since $h(m)\geq 1$), and that $K_{\mathcal{L},0}={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{L}}(m,0)$ by the normal form theorem for amalgamated free products again. Hence for every $m\in\mathbb{Z}$ there is an isomorphism $\phi_{W}^{(m)}:K_{\mathcal{L}}(m,0)\rightarrow K_{\mathcal{L}}(m+1,0)$ sending $J\mapsto J$, $x_{m,0}\mapsto x_{m+1,0}$, and $z_{m,0}\mapsto z_{m+1,0}$, and these extend to the desired automorphism $\phi_{W}:K_{\mathcal{L},0}\rightarrow K_{\mathcal{L},0}$. ∎
We can now define $G_{\mathcal{L}}$ as the iterated HNN extensions of $K_{\mathcal{L}}$ by the automorphisms $\phi_{S},\phi_{T}$, and $\phi_{W}$ from Lemma 4.5, using variables $S,T$, and $W$ respectively. This means that
$G_{\mathcal{L}}=\langle K_{\mathcal{L}},S,T,W:$ $\ SJS^{-1}=TJT^{-1}=WJW^{-1}=J,$
$Sx_{mi}S^{-1}=x_{m,i+1},Tz_{mi}T^{-1}=z_{m,i+1},m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1},$
$Wx_{m,0}W^{-1}=x_{m+1,0},Wz_{m,0}W^{-1}=z_{m+1,0},m\in\mathbb{Z}\rangle.$
Unlike $K_{\mathcal{L}}$, this group is finitely generated:
Lemma 4.6.
$G_{\mathcal{L}}$ is generated by $\{J,S,T,W,x_{00},z_{00}\}$.
Proof.
For any $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$, we have $x_{mi}=S^{i}W^{m}x_{00}W^{-m}S^{-i}$ and $z_{mi}=T^{i}W^{m}z_{00}W^{-m}T^{-i}$ in $G_{\mathcal{L}}$. ∎
We use this generating set frequently going forward. Hence for notational convenience, we let $X:=x_{00}$, $Z:=z_{00}$, and $\mathcal{X}_{G}:=\{J,S,T,W,X,Z\}$. Also, we let $X_{mi}$ and $Z_{mi}$ denote the words
$X_{mi}:=S^{i}W^{m}XW^{-m}S^{-i}\text{ and }Z_{mi}:=T^{i}W^{m}ZW^{-m}T^{-i},m,i\in\mathbb{Z}.$
The words $X_{mi}$ and $Z_{mi}$ both have length $2|m|+2|i|+1$.
If we restate the presentation of $G_{\mathcal{L}}$ with respect to this generating set, we see that $G_{\mathcal{L}}$ is the group generated by $\mathcal{X}_{G}$, subject to the following relations:
  1. (G0)
    $J^{2}=X^{2}=Z^{2}=1$.
  2. (G1)
    $[S,J]=[T,J]=[W,J]=[X,J]=[Z,J]=1$, i.e., $J$ is a central element.
  3. (G2)
    For every $m,i\in\mathbb{Z}$,
    $X_{mi},Z_{mi}]=\begin{cases}J&i+1\neq 0\text{ mod }h(m)+1\\ 1&i+1=0\text{ mod }h(m)+1\end{cases}.$
  4. (G3)
    For all $m\in\mathbb{Z}$ and $i\neq j$ mod $h(m)+1$,
    $X_{mi},Z_{mj}]=[X_{mi},X_{mj}]=[Z_{mi},Z_{mj}]=1.$
  5. (G4)
    For every $m\in\mathbb{N}$ with $h(m)<+\infty$,
    $X_{m,h(m)+1}=X_{m,0}\text{ and }Z_{m,h(m)+1}=Z_{m,0}.$
In the next stage of our construction, we want to use Theorem 2.4 to embed $G_{\mathcal{L}}$ into a finitely presented group with polynomial isoperimetric function. For this, we need:
Proposition 4.7.
The word problem for $G_{\mathcal{L}}$ with respect to $\mathcal{X}_{G}$ is solvable in deterministic polynomial time.
Proof.
First consider the larger generating set $\widehat{\mathcal{X}}_{G}:=\widetilde{\mathcal{X}}_{K}\cup\{S,T,W\}$ where $\widetilde{\mathcal{X}}_{K}=\left\{J,x_{mi},z_{mi},m\in\mathbb{Z},-\lfloor\tfrac{h(m)+3}{2}\rfloor<i<\lfloor\tfrac{h(m)+2}{2}\rfloor\right\}$ as in Proposition 4.2. The word problem with respect to $\widehat{\mathcal{X}}_{G}$ is solvable using Britton’s Lemma as follows:
  1. (1)
    Given a word $w$ over $\widehat{\mathcal{X}}_{G}\cup\widehat{\mathcal{X}}_{G}^{-1}$, write
    $w=g_{0}t_{1}^{e_{1}}g_{1}t_{2}^{e_{2}}\cdots t_{n}^{e_{n}}g_{n},$
    where $e_{1},\ldots,e_{n}\in\{\pm 1\}$, $t_{1},\ldots,t_{n}\in\{S,T,W\}$, and $g_{0},\ldots,g_{n}$ are words over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$.
  2. (2)
    For every $0\leq i\leq n$, compute the normal form $\eta(g_{i})$ of $g_{i}$ from Proposition 4.2.
  3. (3)
    If $n=0$, then $w=1$ in $G_{\mathcal{L}}$ if and only if $\eta(g_{0})=1$, and we can stop.
  4. (4)
    Otherwise, use Lemma 4.5 to check whether there is an index $1\leq j<n$ such that $e_{j}=-e_{j+1}$, $t_{j}=t_{j+1}=S$, and $g_{j}\in K_{\mathcal{L},x}$. If such a $j$ exists, then let $g_{j}^{\prime}=J^{c}x_{m_{1},i_{1}^{\prime}}\cdots x_{m_{k},i_{k}^{\prime}}$ where $\eta(g_{j})=J^{c}x_{m_{1},i_{1}}\cdots x_{m_{k},i_{k}}$ and $\lfloor\tfrac{h(m_{l})+3}{2}\rfloor<i_{l}^{\prime}<\lfloor\tfrac{h(m_{l})+2}{2}\rfloor$ is the representative of $i_{l}+e_{j}]\in\mathbb{Z}_{h(m_{l})+1}$ (computed as in Remark 4.4). Return to step (1) with
    $w=g_{0}t_{1}^{e_{1}}\cdots t_{j-1}^{e_{j-1}}(g_{j-1}g_{j}^{\prime}g_{j+1})t_{j+2}^{e_{j+2}}g_{j+2}\cdots t_{n}^{e_{n}}g_{n}.$
  5. (5)
    If no such $j$ exists, perform step (4) with $S,K_{\mathcal{L},x}$ replaced by $T,K_{\mathcal{L},z}$ and $W,K_{\mathcal{L},0}$, with $g_{j}^{\prime}$ adjusted to match the automorphisms $\phi_{T}$ and $\phi_{W}$ from Lemma 4.5 respectively.
  6. (6)
    If no such $j$ exists for $S,T$ and $W$, then $w\neq 1$ in $G_{\mathcal{L}}$ by Britton’s Lemma, and we can stop.
Define the bit-length and maximum index of a word $w$ over $\widehat{\mathcal{X}}_{G}\cup\widehat{\mathcal{X}}_{G}^{-1}$ analogously to words over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$. Specifically, if if $w$ is written as in step (1), then define the maximum index of $w$ to be $\max_{i}I_{i}$, where $I_{i}$ is the maximum index of $g_{i}$. Let $N$, $B$, and $I$ be the length, bit-length, and maximum index of the initial input to the above algorithm. Since $n$ decreases by two after each iteration, and $n$ is bounded by $N$, the algorithm is guaranteed to halt after at most $\lceil N/2\rceil$ iterations. In steps (4) and (5), the length of $g_{j}^{\prime}$ is equal to the length of $\eta(g_{j})$, which is at most one more than the length of $g_{j}$. The generators $t_{j}$ and $t_{j+1}$ are removed from $w$ in this step, and $w$ does not change in any other step, so the length of $w$ also decreases in each iteration. The maximum index can increase by one in steps (4) and (5) when the automorphisms $\phi_{S}$ and $\phi_{T}$ are applied, so the maximum index throughout is bounded by $I+\lceil N/2\rceil$. Applying automorphisms $\phi_{S}$, $\phi_{T}$, and $\phi_{W}$ in steps (4) and (5) can also increase the bit-length of each generator by one (note that the bit-length of the representative of $i]\in\mathbb{Z}_{h(m)+1}$ is at most the bit-length of $i$), so the overall bit-length of $w$ can go up by at most $N$ in each step. Thus the bit-length of $w$ is bounded by $B+N\cdot\lceil N/2\rceil$ throughout.
By Proposition 4.2, each normal form $\eta(g_{j})$ in step (2) can be computed in time polynomial in the bit-length of $w$, and we compute at most $N$ normal forms in this step. In steps (4) and (5), each representative $i_{\ell}^{\prime}$ can be computed in time polynomial in the bit-length and maximum index of $w$, and we have to compute at most $N$ representatives. The other steps can be done in time polynomial in the bit-length of $w$, so we conclude that the entire algorithm takes time polynomial in $N$, $B$, and $I$.
Since $I$ is not polynomial in $N$ and $B$ in general, the above algorithm doesn’t run in polynomial time in $N$ or $B$. However, if the initial input is a word in $\mathcal{X}_{G}\cup\mathcal{X}_{G}^{-1}$, then $I=0$, and $B=CN$ where $C$ is the number of bits needed to encode each element of $\mathcal{X}_{G}\cup\mathcal{X}_{G}^{-1}$. So the above algorithm takes time polynomial in $N$ for words from $\mathcal{X}_{G}\cup\mathcal{X}_{G}^{-1}$. ∎
We can now embed $G_{\mathcal{L}}$ into a finitely presented group:
Proposition 4.8.
There is a finitely presented group $H_{\mathcal{L}}=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$, a polynomial $d_{\mathcal{L}}$, and an embedding $\iota:\mathcal{F}_{\mathcal{X}_{G}}\hookrightarrow\mathcal{F}_{\mathcal{X}_{H}}$ such that
  1. (a)
    $z^{2}\in\mathcal{R}_{H}$ for all $z\in\mathcal{X}_{H}$ (so all the generators of $H_{\mathcal{L}}$ are involutions),
  2. (b)
    $\iota(J)$, $\iota(X)$, and $\iota(Z)$ are elements of $\mathcal{X}_{H}$, while $\iota(S)$, $\iota(T)$, and $\iota(W)$ have length two in $H_{\mathcal{L}}$, and
  3. (c)
[…]
    $\iota(w)=z_{1}r_{1}z_{1}^{-1}z_{2}r_{2}z_{2}^{-1}\cdots z_{k}r_{k}z_{k}^{-1}$
    in $\mathcal{F}_{\mathcal{X}_{H}}$.
Item (c) implies that $\iota$ embeds $G_{\mathcal{L}}$ into $H_{\mathcal{L}}$.
Proof.
By Proposition 4.7 and Theorem 2.4, $G_{\mathcal{L}}$ can be embedded in a finitely presented group $H_{0}=\langle\mathcal{X}_{0}:\mathcal{R}_{0}\rangle$ with polynomial isoperimetric function $d_{H_{0}}$. Let $\mathcal{X}_{1}=\mathcal{X}_{0}\setminus\{J,X,Z\}$, and $\mathcal{X}_{H}=\{J,X,Z\}\cup\{s_{y},t_{y}:y\in\mathcal{X}_{1}\}$. Define $f:\mathcal{F}_{\mathcal{X}_{0}}\to\mathcal{F}_{\mathcal{X}_{H}}$ by $f(y)=y$ for $y\in\{J,X,Z\}$ and $f(y)=s_{y}t_{y}$ for $y\in\mathcal{X}_{1}$, so in other words $f(w)$ is the word $w$ with $y$ replaced by $s_{y}t_{y}$ for all $y\in\mathcal{X}_{1}$. Finally, let $\mathcal{R}_{H}=\{z^{2}:z\in\mathcal{X}_{H}\}\cup\{f(r):r\in\mathcal{R}_{0}\}$. Then
$\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle=\langle H_{0},s_{y},t_{y}\text{ for }y\in\mathcal{X}_{1}:s_{y}^{2}=t_{y}^{2}=1\text{ and }y=s_{y}t_{y}\text{ for all }y\in\mathcal{X}_{1}\rangle,$
and thus $H_{0}$ is a subgroup of $H:=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$ by repeatedly applying Lemma 2.1. The inclusion $\iota:G_{0}\hookrightarrow H_{0}\hookrightarrow H$ satisfies parts (a) and (b) of the proposition. Let $\ell$ be the length of the longest relation in $\mathcal{R}_{0}$. If $w\in\mathcal{F}_{\mathcal{X}_{G}}$ is trivial in $G_{\mathcal{L}}$, then $w$ is also trivial in $H_{0}$. Hence by Lemma 2.3, there are $r_{1},\ldots,r_{k}\in\mathcal{R}_{0}\cup\mathcal{R}_{0}^{-1}$ and $z_{1},\ldots,z_{k}\in\mathcal{F}_{\mathcal{X}_{0}}$ such that
$w=z_{1}r_{1}z_{1}^{-1}z_{2}r_{2}z_{2}^{-1}\cdots z_{k}r_{k}z_{k}^{-1},$
where $k\leq d_{0}(|w|)$ and $|z_{i}|\leq\ell k+\ell+|w|$ for all $1\leq i\leq k$. Thus $\iota(w)=f(z_{1})f(r_{1})f(z_{1})^{-1}\cdots f(z_{k})f(r_{k})f(z_{k})^{-1}$. Since $d_{H_{0}}$ is polynomial, $f(r_{i})\in\mathcal{R}_{H}\cup\mathcal{R}_{H}^{-1}$ for all $1\leq i\leq k$, and $|w|\leq|f(w)|\leq 2|w|$ for all $w\in\mathcal{F}_{\mathcal{X}_{0}}$, part (c) is true with e.g. $d_{\mathcal{L}}(n):=2\ell d_{H_{0}}(n)+2\ell+2n$. ∎
4.2. Construction of $\mathcal{A}_{\mathcal{L}}(m)$
For this section, we fix a finitely presented group $H_{\mathcal{L}}=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$ and a polynomial $d_{\mathcal{L}}$ as in Proposition 4.8, and identify $G_{\mathcal{L}}$ with its image in $H_{\mathcal{L}}$ (so we don’t mention the embedding $\iota$ explicitly). In particular, if $w\in\mathcal{F}_{\mathcal{X}_{G}}$, then $|w|$ will denote the length of $w$ in $H_{\mathcal{L}}$ rather than $G_{\mathcal{L}}$ unless otherwise noted. The variables $J$, $X$, and $Z$ will continue to denote generators of $H_{\mathcal{L}}$, while $S$, $T$ and $W$ will be length two words in $H_{\mathcal{L}}$. We also continue with the notation $X_{mi}=S^{i}W^{m}XW^{-m}S^{-i}$ and $Z_{mi}=T^{i}W^{m}ZW^{-m}T^{-i}$. Now that we are working with algebras, we let $[A,B]$ denote the algebra commutator ABBAAB-BA instead of the group commutator.
To define $\mathcal{A}_{\mathcal{L}}(m)$, consider the set of variables
$\mathcal{X}_{0}:=\{U_{1},U_{2},\widetilde{X},\widetilde{Z},O_{P},O_{Q}\},$
and let
$U:=U_{1}U_{2},\quad P:=\frac{1-O_{P}}{2},\text{ and }Q:=\frac{1-O_{Q}}{2}$
in $\mathbb{C}^{*}\langle\mathcal{X}_{0}\rangle$. For every $m\in\mathbb{N}$, let $\mathcal{A}_{\mathcal{L}}(m)$ be the finitely presented *-algebra generated by $\mathcal{X}_{H}\cup\mathcal{X}_{0}$, subject to the relations
  1. (R0)
    $x^{*}x=xx^{*}=x^{2}=1$ for all $x\in\mathcal{X}_{H}\cup\mathcal{X}_{0}$,
  2. (R1)
    $r=1$ for all $r\in\mathcal{R}_{H}$, i.e. all the relations for $H_{\mathcal{L}}$ hold,
  3. (R2)
    [U,X]=[U,Z]=[U,S]=[U,T]=[U,J]=[Q,X]=[Q,Z]=[Q,S]=[Q,T]=[Q,J]=0[U,X]=[U,Z]=[U,S]=[U,T]=[U,J]=[Q,X]=[Q,Z]=[Q,S]=[Q,T]=[Q,J]=0, i.e. $U$ and $Q$ commute with $X$,$Z$,$S$,$T$, and $J$,
  4. (R3)
    $\widetilde{X},Q]=\widetilde{X}Q-X_{m,0}Q=0$ and $\widetilde{Z},Q]=\widetilde{Z}Q-Z_{m,0}Q=0$,
  5. (R4)
    $U\widetilde{X}U^{*}-S\widetilde{X}S^{*}=U\widetilde{Z}U^{*}-T\widetilde{Z}T^{*}=0$,
  6. (R5)
    $[P,Q]=0$, i.e. $P$ commutes with $Q$, and
  7. (R6)
    $(P+\widetilde{X}P\widetilde{X}-UPU^{*})(1+J\widetilde{X}\widetilde{Z}\widetilde{X}\widetilde{Z})=0$.
Note that relations (R0) imply that all the generators are order-two unitaries, so $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}/\langle\mathcal{R}_{m}\rangle$, where $\mathcal{R}_{m}$ denotes the relations in (R1)-(R6). Since $O_{P}$ and $O_{Q}$ have order two, $P$ and $Q$ are projections in $\mathcal{A}_{\mathcal{L}}(m)$. The point of the generators $O_{P}$ and $O_{Q}$ is to add the projections $P$ and $Q$ while making $\mathcal{A}_{\mathcal{L}}(m)$ a quotient of a free product of $\mathbb{Z}_{2}$’s. Similarly, the point of $U_{1}$ and $U_{2}$ is to add the unitary $U$ without putting any restriction on its order. For the intuition behind the remaining relations, see the representation constructed in the proof of Lemma 4.10 below. Before proceeding to the lemma, we record some facts about the presentation for later use:
Remark 4.9.
  1. (1)
    If $\mathcal{A}=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}$, then $\lVert r\rVert_{\mathcal{A}}\leq 6$ for all $r\in\mathcal{R}_{m}$.
  2. (2)
    The number of relations in $\mathcal{R}_{m}$ is independent of $m$.
  3. (3)
    The relations $r$ in (R3) satisfy $\lVert r\rVert_{1,1}\leq 2m+3$, and the other relations in $\mathcal{R}_{m}$ do not depend on $m$.
Lemma 4.10.
If $m\in\mathcal{L}$ then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(PQ)>0$.
Proof.
[…]
$\widetilde{\pi}(x)=\begin{pmatrix}\pi(x)\otimes Id_{\mathcal{H}_{1}}&\quad\\ \quad&\pi(x)\otimes Id_{\mathcal{H}_{1}}\end{pmatrix}\text{ for all }x\in\mathcal{X}_{H},$
$\widetilde{\pi}(U_{1})=\begin{pmatrix}\quad&Id_{\mathcal{H}_{0}}\otimes\zeta\\ Id_{\mathcal{H}_{0}}\otimes\zeta^{-1}\end{pmatrix},\quad\widetilde{\pi}(U_{2})=\begin{pmatrix}\quad&Id_{\widetilde{\mathcal{H}}}\\ Id_{\widetilde{\mathcal{H}}}\end{pmatrix}$,
$\widetilde{\pi}(\widetilde{X})=\begin{pmatrix}\sum\limits_{i=0}^{n}\pi(X_{mi})\otimes E_{-i,-i}&\quad\\ \quad&\sum\limits_{i=0}^{n}\pi(X_{mi})\otimes E_{i,i}\end{pmatrix}$,
$\widetilde{\pi}(\widetilde{Z})=\begin{pmatrix}\sum\limits_{i=0}^{n}\pi(Z_{mi})\otimes E_{-i,-i}&\quad\\ \quad&\sum\limits_{i=0}^{n}\pi(Z_{mi})\otimes E_{i,i}\end{pmatrix}$,
$\widetilde{\pi}(Q)=\begin{pmatrix}Id_{\mathcal{H}_{0}}\otimes E_{0,0}&\quad\\ \quad&Id_{\mathcal{H}_{0}}\otimes E_{0,0}\end{pmatrix},\text{ and }$
$\widetilde{\pi}(P)=\begin{pmatrix}\sum\limits_{i=0}^{n}\pi(P_{i})\otimes E_{-i,-i}&\quad\\ \quad&\sum\limits_{i=0}^{n}\pi(P_{i})\otimes E_{i,i}\end{pmatrix}$,
where in the last equation $P_{i}:=(\frac{1-J}{2})(\frac{1-Z_{m,i}}{2})(\frac{1-Z_{m,i+1}}{2})\cdots(\frac{1-Z_{m,n-1}}{2})$ for all $0\leq i\leq n-1$, and $P_{n}:=\frac{1-J}{2}$. Note that
$\widetilde{\pi}(U)=\widetilde{\pi}(U_{1})\widetilde{\pi}(U_{2})=\begin{pmatrix}Id_{\mathcal{H}_{0}}\otimes\zeta&\quad\\ \quad&Id_{\mathcal{H}_{0}}\otimes\zeta^{-1}\end{pmatrix}$.
It is not hard to check that $\widetilde{\pi}(r)=0$ for all relations $r$ in (R0)-(R6). Indeed, it is clear that $\widetilde{\pi}(x)$ is a unitary squaring to $1$ for $x\in\mathcal{X}_{H}\cup\{U_{1},U_{2},\widetilde{X},\widetilde{Z}\}$, and $\widetilde{\pi}(Q)$ is a projection. Since $Z_{m,0},\cdots,Z_{m,n-1}$ and $J$ commute in $\mathbb{C}H_{\mathcal{L}}$, $\widetilde{\pi}(P)$ is a projection as well. So $\widetilde{\pi}(r)=0$ for all $r$ in (R0). We also have $\widetilde{\pi}(r)=0$ for $r$ in (R1), (R2), and (R5) by inspection. For (R3),
$\widetilde{\pi}(\widetilde{X})\widetilde{\pi}(Q)=\widetilde{\pi}(Q)\widetilde{\pi}(\widetilde{X})=\widetilde{\pi}(X_{m,0})\widetilde{\pi}(Q)=\begin{pmatrix}\pi(X_{m,0})\otimes E_{0,0}&\quad\\ \quad&\pi(X_{m,0})\otimes E_{0,0}\end{pmatrix}$,
and similarly $\widetilde{\pi}(\widetilde{Z})\widetilde{\pi}(Q)=\widetilde{\pi}(Q)\widetilde{\pi}(\widetilde{Z})=\widetilde{\pi}(Z_{m,0})\widetilde{\pi}(Q)$. For (R4), recall from the presentation of $G_{\mathcal{L}}$ that $\pi(X_{m,0})=\pi(X_{m,n+1})$ and $\pi(S)\pi(X_{m,i})\pi(S^{*})=\pi(X_{m,i+1})$ for all $i\in\mathbb{Z}$. Since $\zeta E_{i,i}\zeta^{-1}=E_{i+1,i+1}$ for all $i\in\mathbb{Z}_{n+1}$,
$\widetilde{\pi}(S)\widetilde{\pi}(\widetilde{X})\widetilde{\pi}(S^{*})$ $=\begin{pmatrix}\sum\limits^{n}_{i=0}\pi(X_{m,i+1})\otimes E_{-i,-i}&\quad\\ \quad&\sum\limits^{n}_{i=0}\pi(X_{m,i+1})\otimes E_{i,i}\end{pmatrix}$
$=\begin{pmatrix}\sum\limits^{n}_{i=0}\pi(X_{m,i})\otimes E_{-(i-1),-(i-1)}&\quad\\ \quad&\sum\limits^{n}_{i=0}\pi(X_{m,i})\otimes E_{i-1,i-1}\end{pmatrix}$
$=\widetilde{\pi}(U)\widetilde{\pi}(\widetilde{X})\widetilde{\pi}(U^{*}).$
Similarly $\widetilde{\pi}(T)\widetilde{\pi}(\widetilde{Z})\widetilde{\pi}(T^{*})=\widetilde{\pi}(U)\widetilde{\pi}(\widetilde{Z})\widetilde{\pi}(U^{*})$. This just leaves (R6). For this case,
$\widetilde{\pi}\big((P+\widetilde{X}P\widetilde{X}-UPU^{*})(1+J\widetilde{X}\widetilde{Z}\widetilde{X}\widetilde{Z})\big)=\begin{pmatrix}A_{1}&\quad\\ \quad&A_{2}\end{pmatrix}\text{ where}$
$A_{1}=\sum_{i=0}^{n}\pi\big((P_{i}+X_{mi}P_{i}X_{mi}-P_{i+1})(1+JX_{mi}Z_{mi}X_{mi}Z_{mi})\big)\otimes E_{-i,-i},$
$A_{2}=\sum_{i=0}^{n}\pi\big((P_{i}+X_{mi}P_{i}X_{mi}-P_{i+1})(1+JX_{mi}Z_{mi}X_{mi}Z_{mi})\big)\otimes E_{i,i},$
and $P_{n+1}:=P_{0}$. Recall from the presentation of $G_{\mathcal{L}}$ that $J$ is a unitary of order 2 and is central. Hence in $\mathbb{C}H_{\mathcal{L}}$, $\tfrac{1-J}{2}$ and $\tfrac{1+J}{2}$ are projections, and $(\tfrac{1+J}{2})P_{i}=0$ for $0\leq i\leq n$. Also for all $0\leq i\leq n-1$, $X_{mi}Z_{mi}X_{mi}Z_{mi}=J$, $X_{mi}$ commutes with $P_{i+1}$, and $P_{i}=(\tfrac{1-Z_{mi}}{2})P_{i+1}$. Hence $1+JX_{mi}Z_{mi}X_{mi}Z_{mi}=2$, and
$P_{i}+X_{mi}P_{i}X_{mi}-P_{i+1}=(\tfrac{1-Z_{mi}}{2})P_{i+1}+(\tfrac{1-JZ_{mi}}{2})P_{i+1}-P_{i+1}=-Z_{mi}(\tfrac{1+J}{2})P_{i+1}=0.$
For $i=n$, $X_{mn}Z_{mn}X_{mn}Z_{mn}=1$, so
$(P_{n}+X_{mn}P_{n}X_{mn}-P_{0})(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})=(2-P_{0})\tfrac{(1-J)(1+J)}{2}=0.$
It follows that $\pi\big((P_{i}+X_{mi}P_{i}X_{mi}-P_{i+1})(1+JX_{mi}Z_{mi}X_{mi}Z_{mi})\big)=0$ for all $0\leq i\leq n$, and hence $A_{1}=A_{2}=0$. We conclude that $\widetilde{\pi}(r)=0$ for all $r$ in (R0)-(R6). Hence $\widetilde{\pi}$ induces a representation of $\mathcal{A}_{\mathcal{L}}(m)$ on $\widetilde{\mathcal{H}}\oplus\widetilde{\mathcal{H}}$.
The image of $\widetilde{\pi}$ is contained in $M_{2}\big(\pi(\mathbb{C}H_{\mathcal{L}})\otimes\mathcal{B}(\mathcal{H}_{1})\big)$ This algebra is isomorphic to $\pi(\mathbb{C}H_{\mathcal{L}})\otimes M_{2(n+1)}(\mathbb{C})$, which has a tracial state $\tau:=\tau_{0}\otimes\tau^{\prime}$ where $\tau^{\prime}$ is the unique tracial state on $M_{2(n+1)}(\mathbb{C})$. Finally,
$\tau\big(\widetilde{\pi}(PQ)\big)$ $=\tau\begin{pmatrix}\pi(P_{0})\otimes E_{0,0}&\quad\\ \quad&\pi(P_{0})\otimes E_{0,0}\end{pmatrix}=\tfrac{1}{(n+1)}\tau_{0}(P_{0})=\tfrac{1}{2^{n+1}(n+1)}>0,$
so the lemma follows. ∎
We now look at what happens when $m\not\in\mathcal{L}$. For all $n\geq 0$, let
$\widetilde{P}_{n}:=QU^{n}PU^{-n}Q\text{ and }\widetilde{X}_{n}:=U^{n}\widetilde{X}U^{-n}$
in $\mathbb{C}^{*}\langle\mathcal{X}_{H}\cup\mathcal{X}_{0}\rangle$. Note that $\widetilde{X}_{n}^{2}=1$ in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}$.
Proposition 4.11.
There are constants $C$ and $k$ such that for any $m\in\mathbb{N}$ and $0\leq n<h(m)$, $\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}$ is trivial in $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}/\langle\mathcal{R}_{m}\rangle$, and has an $\mathcal{R}_{m}$-decomposition in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}$ of size $\leq C\big((n+1)m\big)^{k}$.
Although we won’t need this, the constants $C$ and $k$ can be computed from a description of the Turing machine $M_{\mathcal{L}}$ that recognizes $\mathcal{L}$.
Proof.
For the proof, say that $\alpha(n)=O(\beta(n))$ for functions $\alpha,\beta$ if there is a constant $C>0$ such that $\alpha(n)\leq C\beta(n)+C$ for all $n\geq 0$. Recall from Remark 4.9 that $\lVert r\rVert_{\mathcal{A}}\leq 6$ for all $r\in\mathcal{R}_{m}$ when $\mathcal{A}=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}$. Let
$f=QU^{n}(P+\widetilde{X}P\widetilde{X}-UPU^{*})U^{-n}\text{ and }g=U^{n}(1+J\widetilde{X}\widetilde{Z}\widetilde{X}\widetilde{Z})U^{-n}Q.$
By relation (R6), $fg$ is trivial in $\mathcal{A}_{\mathcal{L}}(m)$ and has an $\mathcal{R}_{m}$-decomposition in $\mathcal{A}$ of size $6n+4$. To prove the proposition, we show that
$fg=2\left(\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}\right)$
in $\mathcal{A}_{\mathcal{L}}(m)$, while keeping track of the size using Lemma 3.3. For the first step, since $U$ and $S$ commute and $U\widetilde{X}U^{*}=S\widetilde{X}S^{*}$, we see that $\widetilde{X}_{n}=S^{n}\widetilde{X}S^{-n}$ in $\mathcal{A}_{\mathcal{L}}(m)$. Similarly $U^{n}\widetilde{Z}U^{-n}=T^{n}\widetilde{Z}T^{-n}$ and $U$ commutes with $J$, so
$g=(1+JS^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n}\cdot S^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n})Q.$
Replacing $U^{n}\widetilde{X}U^{-n}$ and $U^{n}\widetilde{Z}U^{-n}$ with $S^{n}\widetilde{X}S^{-n}$ and $T^{n}\widetilde{Z}T^{-n}$ requires $O(n^{2})$ applications of the defining relations, and the intermediate polynomials in this calculation have degree $O(n)$, so by Lemma 3.3,
$g-(1+JS^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n}\cdot S^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n})Q$
has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3})$.
Next, $\widetilde{X}_{n}Q=Q\widetilde{X}_{n}=S^{n}\widetilde{X}QS^{-n}=QX_{mn}$ and $U^{n}\widetilde{Z}U^{-n}Q=QZ_{mn}$ for all $n\geq 0$ by relations (R2) and (R3), so
$(1+JS^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n}\cdot S^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n})Q=Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn}).$
This identity requires using the relations $O(n)$ times, and $X_{mn}$ and $Z_{mn}$ have degree $O(m+n)$, so by Lemma 3.3 again,
$(1+JS^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n}\cdot S^{n}\widetilde{X}S^{-n}\cdot T^{n}\widetilde{Z}T^{-n})Q-Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})$
has an $\mathcal{R}_{m}$-decomposition of size $O(n^{2}+nm)$.
Now suppose that $0\leq n<h(m)$. From the presentation of $G_{\mathcal{L}}$, we see that $X_{mn}Z_{mn}X_{mn}Z_{mn}=J$, so $Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})=Q(1+J^{2})=2Q$. The word $JX_{mn}Z_{mn}X_{mn}Z_{mn}$ has length $16m+16n+5$, so by Proposition 4.8, part (c) there are $r_{1},\ldots,r_{k}\in\mathcal{R}_{H}\cup\mathcal{R}_{H}^{-1}$ and $z_{1},\ldots,z_{k}\in\mathcal{F}_{\mathcal{X}_{H}}$ with $JX_{mn}Z_{mn}X_{mn}Z_{mn}=z_{1}r_{1}z_{1}^{-1}\cdots z_{k}r_{k}z_{k}^{-1}$ in $\mathcal{F}_{\mathcal{X}_{H}}$, $k\leq d_{\mathcal{L}}(16m+16n+5)$, and $|z_{i}|\leq d_{\mathcal{L}}(16m+16n+5)$ for all $1\leq i\leq k$. Replacing $r_{i}$ with $1$ from right to left, we see that $Q(1+JX_{mn}Z_{mn}X_{mn}Z_{mn})-2Q$ has an $\mathcal{R}_{m}$-decomposition of size $O(d_{\mathcal{L}}(16m+16n+5)^{2})=O(d_{\mathcal{L}}(m)^{2}+d_{\mathcal{L}}(n)^{2})$. From Lemma 3.3, part (c), we get that fg2fQfg-2fQ has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3}+n^{2}+nm+d_{\mathcal{L}}(m)^{2}+d_{\mathcal{L}}(n)^{2})$.
We now look at $fQ$. Since $Q$ commutes with $\widetilde{X}_{n}$,
$fQ$ $=QU^{n}PU^{-n}Q+QU^{n}\widetilde{X}P\widetilde{X}U^{-n}Q-QU^{n+1}PU^{-(n+1)}Q$
$=\widetilde{P}_{n}+Q\widetilde{X}_{n}\cdot U^{n}PU^{-n}\cdot\widetilde{X}_{n}Q-\widetilde{P}_{n+1}$
$=\widetilde{P}_{n}+\widetilde{X}_{n}\cdot\widetilde{P}_{n}\cdot\widetilde{X}_{n}-\widetilde{P}_{n+1}.$
This identity requires replacing $\widetilde{X}_{n}$ with $S^{n}\widetilde{X}S^{-n}$ again, so
$fQ-(\widetilde{P}_{n}+\widetilde{X}_{n}\cdot\widetilde{P}_{n}\cdot\widetilde{X}_{n}-\widetilde{P}_{n+1})$
has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3})$. Ultimately we conclude that $\widetilde{P}_{n}+\widetilde{X}_{n}\cdot\widetilde{P}_{n}\cdot\widetilde{X}_{n}-\widetilde{P}_{n+1}$ has an $\mathcal{R}_{m}$-decomposition of size $O(n^{3}+n^{2}+nm+d_{\mathcal{L}}(m)^{2}+d_{\mathcal{L}}(n)^{2})$. ∎
Proof of Theorem 4.1.
Let $\mathcal{X}=\mathcal{X}_{H}\cup\mathcal{X}_{0}$, and let $\widetilde{P}_{n}$, $\widetilde{X}_{n}$, and $\mathcal{R}_{m}$ be as above. The relations in (R1) are independent of $m$, and the remaining relations are spelled out explicitly in (R2)-(R6), so we see that the mapping $m\mapsto\mathcal{R}_{m}$ is computable. Parts (a) and (d) of the theorem hold by construction, and part (b) follows from Proposition 4.11. Finally, since $[P,Q]=0$ in $\mathcal{A}_{\mathcal{L}}(m)$, Lemma 4.10 implies that there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ with $\tau(\widetilde{P}_{0}^{2})=\tau(PQ)>0$. ∎
5. Proof of Theorem 1.3
In this section, we prove the following theorem.
Theorem 5.1.
There is an integer $n_{0}$, such that for all $N\geq n_{0}$, the promise problem
Given $\alpha\in\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, determine if $\alpha$ is trace-positive in $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, or if $\pi(\alpha)$ is not trace-positive in $\mathbb{C}\mathbb{Z}_{2}^{*N}$, promised that one of the two is the case.
is coRE-hard, where $\pi:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ is the homomorphism sending $x_{i}$ to the generator of the $i$th factor of $\mathbb{Z}_{2}^{*N}$.
Suppose that $N\geq n_{0}$, and that the homomorphism $\pi$ in Theorem 5.1 factors through a *-algebra $\mathcal{A}$, meaning that there are homomorphisms $\pi_{1}:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathcal{A}$ and $\pi_{2}:\mathcal{A}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ such that $\pi=\pi_{2}\circ\pi_{1}$. Suppose also that we have some way of representing the elements of $\mathcal{A}$ as strings such that the homomorphism $\pi_{1}$ is computable. Then Theorem 5.1 has the following immediate corollary:
Corollary 5.2.
It is coRE-hard to determine whether $\alpha\in\mathcal{A}$ is trace-positive in $\mathcal{A}_{\mathbb{C}}$.
Proof.
Suppose $\alpha\in\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$. If $\alpha$ is trace-positive in $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, then $\pi_{1}(\alpha)$ is trace-positive in $\mathcal{A}_{\mathbb{C}}$. If $\pi(\alpha)$ is not trace-positive in $\mathbb{C}\mathbb{Z}_{2}^{*N}$, then there is a tracial state $\tau$ on $\mathbb{C}\mathbb{Z}_{2}^{*N}$ such that $(\tau\circ\pi_{2})(\pi_{1}(\alpha))=\tau(\pi(\alpha))<0$, so $\pi_{1}(\alpha)$ is not trace-positive in $\mathcal{A}_{\mathbb{C}}$. Since $\pi_{1}$ is computable, the promise problem in Theorem 5.1 reduces to the problem of determining if $\alpha\in\mathcal{A}$ is trace-positive. ∎
Theorem 1.3 is a special case of 5.2 with $\mathcal{A}=\mathbb{Q}^{*}\{x_{1},\ldots,x_{N}\}$, $\mathbb{Q}\mathcal{F}_{N}$, or $\mathbb{Q}\mathbb{Z}_{M}^{*N}$ (for every $M\geq M^{\prime}$, there is a *-homomorphism $\mathbb{Q}\mathbb{Z}_{M}^{*N}\rightarrow\mathbb{Q}\mathbb{Z}_{M^{\prime}}^{*N}$ sending the $i$th generator of $\mathbb{Q}\mathbb{Z}_{M}^{*N}$ to the $i$th generator of $\mathbb{Q}\mathbb{Z}_{M^{\prime}}^{*N}$). The corollary also applies to other algebras such as the algebra of contractions and the algebra of self-adjoint contractions, as well as free products of any of the algebras mentioned.
For the proof of Theorem 5.1, suppose $\mathcal{L}\subseteq\mathbb{N}$ is an RE set, and let $\mathcal{X}$, $\widetilde{P}_{n}$ and $\widetilde{X}_{n}$, $n\geq 0$, and $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$, $m\geq 1$, be as in Theorem 4.1. Before proceeding to the proof, we need the following propositions.
Proposition 5.3.
There are positive integers $\Lambda,k$ such that if $m\notin\mathcal{L}$ and $\tau$ is a tracial $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for some $\epsilon\geq 0$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Lambda m^{k}\sqrt{\epsilon}$.
Proof.
Let $C$ and $k$ be the constants from Theorem 4.1, part (b). Suppose $m\notin\mathcal{L}$, $\epsilon\geq 0$, and $\tau$ is a tracial $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Theorem 4.1, part (a), $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. Since $\tau$ is tracial, $\lVert\widetilde{P}_{n}\rVert_{\tau}=\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}$ for all $n$. Also $\widetilde{P}_{n}$ and $\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}$ are hermitian squares in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, so $\tau(\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n})=\tau(\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n})\geq 0$. Hence
$\lVert\widetilde{P}_{n}\rVert_{\tau}^{2}=\tau(\widetilde{P}_{n}^{2})$ $=\tfrac{1}{2}\Big(\tau(\widetilde{P}_{n}^{2})+\tau\big((\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n})^{2}\big)\Big)$
$\leq\tfrac{1}{2}\Big(\tau(\widetilde{P}_{n}^{2})+\tau\big((\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n})^{2}\big)+\tau(\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n})+\tau(\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n})\Big)$
$=\tfrac{1}{2}\tau\big((\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n})^{2}\big)=\tfrac{1}{2}\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}^{2}.$
By Proposition 3.4 and Theorem 4.1, part (b), $\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}\rVert_{\tau}\leq C((n+1)m)^{k}\sqrt{\epsilon}$ for all $n\geq 0$. It follows that
$\lVert\widetilde{P}_{n}\rVert_{\tau}$ $\leq\tfrac{1}{\sqrt{2}}\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}$
$\leq\tfrac{1}{\sqrt{2}}(\lVert\widetilde{P}_{n+1}\rVert_{\tau}+\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}\rVert_{\tau})$
$\leq\tfrac{1}{\sqrt{2}}\lVert\widetilde{P}_{n+1}\rVert_{\tau}+\tfrac{1}{\sqrt{2}}C((n+1)m)^{k}\sqrt{\epsilon}.$
By induction, we obtain
$\lVert\widetilde{P}_{0}\rVert_{\tau}\leq(\tfrac{1}{\sqrt{2}})^{t}\lVert\widetilde{P}_{t}\rVert_{\tau}+C\sum_{n=1}^{t}\frac{n^{k}}{2^{n/2}}m^{k}\sqrt{\epsilon}$
for all $t\geq 1$. Since $\widetilde{P}_{t}$ is a contraction in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ by Theorem 4.1, part (a), $\lVert\widetilde{P}_{t}\rVert_{\tau}\leq 1$ for all $t\geq 1$. The series $\sum_{n=1}^{\infty}\frac{n^{k}}{2^{n/2}}$ converges, so taking $t\rightarrow+\infty$, we conclude that
$\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Lambda m^{k}\sqrt{\epsilon}$
for any integer $\Lambda\geq C\sum_{n=1}^{\infty}\frac{n^{k}}{2^{n/2}}$. ∎
Proposition 5.4.
Let $\mathcal{W}$ be the mapping from Lemma 3.11, and let
$\mathcal{W}_{m}:=\left(\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\right)\cup\mathcal{W}(\widetilde{P}_{0})\cup\mathcal{R}_{m}$
for $m\in\mathbb{N}$. There are positive integers $\widetilde{\Lambda},k^{\prime}$ such that if $m\notin\mathcal{L}$ and $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\widetilde{\Lambda}m^{k^{\prime}}\sqrt{\epsilon}$.
Proof.
Let $\Lambda,k$ be integers from Proposition 5.3, and let $D$ be the integer from Theorem 4.1, part (d). Suppose $m\notin\mathcal{L}$, $\epsilon\geq 0$, and $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle$. By Lemma 3.11, there is a tracial state $\widehat{\tau}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that
$\big|\lVert r\rVert_{\widehat{\tau}}-\lVert r\rVert_{\tau}\big|\leq 2\lVert r\rVert_{1,1}\sqrt{\epsilon}\leq 2Dm\sqrt{\epsilon}$
for all $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$. Since $\mathcal{R}_{m}\subseteq\mathcal{W}_{m}$, $\tau$ is an $(\epsilon,\mathcal{R}_{m})$-state. For every $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$, $\lVert r\rVert_{\tau}\leq\sqrt{\epsilon}$, and hence $\lVert r\rVert_{\widehat{\tau}}\leq(2Dm+1)\sqrt{\epsilon}$. We conclude that $\widehat{\tau}$ is a tracial $\big((2Dm+1)^{2}\epsilon,\mathcal{R}_{m}\big)$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Proposition 5.3, $\lVert\widetilde{P}_{0}\rVert_{\widehat{\tau}}\leq\Lambda m^{k}\sqrt{(2Dm+1)^{2}\epsilon}=(2Dm+1)\Lambda m^{k}\sqrt{\epsilon}$. Again by Lemma 3.11, since $\tau(s^{*}s)\leq\epsilon$ for all $s\in\mathcal{W}(\widetilde{P}_{0})$, we have
$\big|\lVert\widetilde{P}_{0}\rVert_{\widehat{\tau}}-\lVert\widetilde{P}_{0}\rVert_{\tau}\big|\leq 2\lVert\widetilde{P}_{0}\rVert_{1,1}\sqrt{\epsilon}=3\sqrt{\epsilon}.$
It follows that
$\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\left((2Dm+1)\Lambda m^{k}+3\right)\sqrt{\epsilon}\leq 6D\Lambda m^{k+1}\sqrt{\epsilon}.$
Taking $\widetilde{\Lambda}:=6D\Lambda$ and $k^{\prime}:=k+1$ completes the proof. ∎
Proof of Theorem 5.1 .
Let $n_{0}:=\lvert\mathcal{X}\rvert$, and suppose we are given an integer $N\geq n_{0}$. Pick an ordering on $\mathcal{X}$ so we can identify $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$ with the subalgebra $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{n_{0}}\rangle$ of $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$. We think of $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ and $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ as subalgebras of $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$ and $\mathbb{C}\mathbb{Z}_{2}^{*N}$ similarly. Let $\widetilde{\Lambda}$, $k^{\prime}$, and $\mathcal{W}_{m}$ be as in Proposition 5.4. For any $m\in\mathbb{N}$, define a *-polynomial
(5.1) $\alpha(m)=\sum_{r\in\mathcal{W}_{m}\cup\mathcal{W}_{m}^{*}}r^{*}r-\frac{1}{\widetilde{\Lambda}^{2}m^{2k^{\prime}}}\widetilde{P}_{0}^{*}\widetilde{P}_{0}\in\mathbb{Q}^{*}\langle\mathcal{X}\rangle\subseteq\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle.$
Then $\alpha(m)$ is self-adjoint and the map $\alpha:\mathbb{N}\rightarrow\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$ is computable by Theorems 4.1 and 3.11.
Let $\pi:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ be the homomorphism sending $x_{i}$ to the $i$th generator of $\mathbb{Z}_{2}^{*N}$. Also let $q=q_{2}\circ q_{1}$, where $q_{1}:\mathbb{Q}\mathbb{Z}_{2}^{*N}\to\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the homomorphism sending $x_{i}\mapsto x_{i}$ for $1\leq i\leq n_{0}$ and $x_{i}\mapsto 1$ for $i>n_{0}$, and $q_{2}:\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\to\mathcal{A}_{\mathcal{L}}(m)$ is the quotient map. By Theorem 4.1, if $m\in\mathcal{L}$ then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$. By Lemma 3.11, $\pi(t)=0$ for all $t\in\big(\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\big)\cup\mathcal{W}(\widetilde{P}_{0})$. Since $q(r)=0$ for all $r\in\mathcal{R}_{m}$, $q(\pi(r^{*}r))=0$ for all $r\in\mathcal{W}_{m}\cup\mathcal{W}_{m}^{*}$, and hence $(\tau\circ q)(\pi(\alpha(m)))=-\frac{1}{\widetilde{\Lambda}^{2}m^{2k^{\prime}}}\tau(\widetilde{P}_{0}^{2})<0$. We conclude that $\pi(\alpha(m))$ is not trace-positive.
Now suppose $m\notin\mathcal{L}$, and that $\tau$ is a tracial state on $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$. Let $\epsilon:=\sum_{r\in\mathcal{W}_{m}\cup\mathcal{W}_{m}^{*}}\tau(r^{*}r)$. Then $\tau$ is a tracial $(\epsilon,\mathcal{W}_{m})$-state, so by Proposition 5.4, $\tau(\widetilde{P}^{*}_{0}\widetilde{P}_{0})=\lVert\widetilde{P}_{0}\rVert^{2}_{\tau}\leq\widetilde{\Lambda}^{2}m^{2k^{\prime}}\epsilon$. Hence $\tau(\alpha(m))\geq 0$. Thus the complement $\overline{\mathcal{L}}$ of $\mathcal{L}$ reduces to the promise problem in the theorem statement. Taking $\mathcal{L}$ to be RE-hard, we see that the promise problem is coRE-hard. ∎
6. Proof of Theorem 1.1
The proof of Theorem 1.1 follows the same approach as the proof of Theorem 1.3. We start with a promise problem version of the theorem.
Theorem 6.1.
Suppose either $N\geq 2$ and $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{3}^{*N}$, or $N\geq 3$ and $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$. Let $\mathcal{P}=\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$, and let $\pi:\mathcal{P}\to\mathcal{C}$ be the homomorphism sending $x_{i}$ to the $i$th generator of $\mathcal{C}$. Then the promise problem
Given $\alpha\in\mathcal{P}\otimes_{\mathbb{Q}}\mathcal{P}$, determine if $\alpha$ is positive in $\mathcal{P}_{\mathbb{C}}\otimes_{\mathbb{C}}\mathcal{P}_{\mathbb{C}}$, or if $(\pi\otimes\pi)(\alpha)$ is not positive in $\mathcal{C}_{\mathbb{C}}\otimes_{\mathbb{C}}\mathcal{C}_{\mathbb{C}}$, promised that one of the two is the case.
is coRE-hard.
To get Theorem 1.1 from Theorem 6.1, suppose that $N$, $\mathcal{C}$, $\mathcal{P}$, and $\pi$ are as in the theorem, and that $\pi\otimes\pi$ factors through a *-algebra $\mathcal{A}$, meaning that there are homomorphisms $\pi_{1}:\mathcal{P}\otimes\mathcal{P}\to\mathcal{A}$ and $\pi_{2}:\mathcal{A}\to\mathcal{C}\otimes\mathcal{C}$ such that $\pi\otimes\pi=\pi_{2}\circ\pi_{1}$. Suppose also that the first homomorphism $\pi_{1}$ in this factorization is computable. Then:
Corollary 6.2.
It is coRE-hard to determine whether $\alpha\in\mathcal{A}$ is positive in $\mathcal{A}_{\mathbb{C}}$.
The proof is the same as the proof of 5.2, with the word “trace” deleted. Theorem 1.1 follows immediately from 6.2 by taking $\mathcal{A}$ to be one of $\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\otimes\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$ or $\mathbb{Q}\mathcal{F}_{N}\otimes\mathbb{Q}\mathcal{F}_{N}$ with $N\geq 2$, or $\mathbb{Q}\mathbb{Z}_{M}^{*N}\otimes\mathbb{Q}\mathbb{Z}_{M}^{*N}$ with $N\geq 3,M\geq 2$ or $N\geq 2,M\geq 3$. As with 5.2, 6.2 also applies to other algebras such as the algebra of contractions and the algebra of self-adjoint contractions, as well as free products of any of the algebras mentioned.
For the proof of Theorem 6.1, we use the same setup as in the previous section: we suppose $\mathcal{L}\subseteq\mathbb{N}$ is an RE set, and let $\mathcal{X},\widetilde{P}_{n},\widetilde{X}_{n}$, $n\geq 0$, and $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}^{*}\langle\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:\mathcal{R}_{m}\rangle$, $m\in\mathbb{N}$, be as in Theorem 4.1. In addition, let $\iota:\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\rightarrow\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}:a\mapsto a\otimes 1$ be the left inclusion.
Proposition 6.3.
There are positive integers $\Gamma,k$ such that if $m\notin\mathcal{L}$, $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, and $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ for some $\epsilon\geq 0$, then $\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Gamma m^{k}\sqrt{\epsilon}$.
Proof.
Suppose $m\notin\mathcal{L}$, $\epsilon\geq 0$, and $\varphi$ is an $(\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that $\tau:=\varphi\circ\iota$ is an $(\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Theorem 4.1, part (a), $\widetilde{P}_{n}$ is a contraction in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, $\widetilde{X}_{n}$ is a monomial in $\mathbb{C}^{*}\langle\mathcal{X}\rangle$ of degree $4n+1$, and $\widetilde{X}_{n}$ is a unitary involution in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, so Lemma 3.7 implies
$\left|\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}-\lVert\widetilde{P}_{n}\rVert_{\tau}\right|=\left|\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\otimes 1\rVert_{\varphi}-\lVert\widetilde{P}_{n}\otimes 1\rVert_{\varphi}\right|\leq(4n+1)\sqrt{\epsilon}$
for all $n\geq 0$. Hence
$\left|\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert^{2}_{\tau}-\lVert\widetilde{P}_{n}\rVert^{2}_{\tau}\right|$ $=\big(\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}+\lVert\widetilde{P}_{n}\rVert_{\tau}\big)\left|\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}-\lVert\widetilde{P}_{n}\rVert_{\tau}\right|$
$\leq\big(2\lVert\widetilde{P}_{n}\rVert_{\tau}+(4n+1)\sqrt{\epsilon}\big)(4n+1)\sqrt{\epsilon},$
so
(6.1) $\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert^{2}_{\tau}\geq\lVert\widetilde{P}_{n}\rVert^{2}_{\tau}-(8n+2)\lVert\widetilde{P}_{n}\rVert_{\tau}\sqrt{\epsilon}-(4n+1)^{2}\epsilon$
for all $n\geq 0$. By Theorem 4.1, part (a) again, $\widetilde{P}_{n}$ is a hermitian square in $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ with $\lVert\widetilde{P}_{n}\rVert_{1}=1$ and $\lVert\widetilde{P}_{n}\rVert_{1,1}=3(n+1)/2\leq 2n+2$ for every $n\geq 0$, so $\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{1,1}\leq\lVert\widetilde{P}_{n}\rVert_{1,1}+\lVert\widetilde{P}_{n}\rVert_{1}(8n+2)\leq 10n+4$. Lemma 3.8 implies
$\varphi\big(\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\otimes 1+\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\otimes 1\big)$ $=2\operatorname{Re}\Big(\varphi\big(\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\otimes 1\big)\Big)$
$\geq-2\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{1,1}\lVert\widetilde{P}_{n}\otimes 1\rVert_{\varphi}\sqrt{\epsilon}$
$\geq-(20n+8)\lVert\widetilde{P}_{n}\rVert_{\tau}\sqrt{\epsilon}.$
By Equation 6.1, we obtain
$\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert^{2}_{\tau}$ $=\lVert\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}^{2}+\lVert\widetilde{P}_{n}\rVert_{\tau}^{2}+\varphi(\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\otimes 1+\widetilde{P}_{n}\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\otimes 1)$
$\geq 2\lVert\widetilde{P}_{n}\rVert_{\tau}^{2}-(28n+10)\lVert\widetilde{P}_{n}\rVert_{\tau}\sqrt{\epsilon}-(4n+1)^{2}\epsilon$
$\geq 2\big(\lVert\widetilde{P}_{n}\rVert_{\tau}-(7n+7)\sqrt{\epsilon}\big)^{2}-(11n+11)^{2}\epsilon,$
where the coefficients here (and below) are chosen for convenience. It follows that
$2\big(\lVert\widetilde{P}_{n}\rVert_{\tau}-(7n+7)\sqrt{\epsilon}\big)^{2}\leq\big(\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}+(11n+11)\sqrt{\epsilon}\big)^{2},$
and hence
$\lVert\widetilde{P}_{n}\rVert_{\tau}$ $\leq\frac{1}{\sqrt{2}}\big(\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}\rVert_{\tau}+(25n+25)\sqrt{\epsilon}\big).$
Let $C$ and $k$ be the constants from Theorem 4.1, part (b). By Lemma 3.9, $\lVert\widetilde{P}_{n}+\widetilde{X}_{n}\widetilde{P}_{n}\widetilde{X}_{n}-\widetilde{P}_{n+1}\rVert_{\tau}\leq C((n+1)m)^{k}\sqrt{\epsilon}$ for all $n\geq 0$. Thus
$\lVert\widetilde{P}_{n}\rVert_{\tau}$ $\leq\tfrac{1}{\sqrt{2}}\big(\lVert\widetilde{P}_{n+1}\rVert_{\tau}+C((n+1)m)^{k}\sqrt{\epsilon}+(25n+25)\sqrt{\epsilon}\big)$
$\leq\tfrac{1}{\sqrt{2}}\lVert\widetilde{P}_{n+1}\rVert_{\tau}+\tfrac{1}{\sqrt{2}}(C+25)((n+1)m)^{k}\sqrt{\epsilon}.$
By induction, we obtain
$\lVert\widetilde{P}_{0}\rVert_{\tau}\leq(\tfrac{1}{\sqrt{2}})^{t}\lVert\widetilde{P}_{t}\rVert_{\tau}+(C+25)\sum_{n=1}^{t}\frac{n^{k}}{2^{n/2}}m^{k}\sqrt{\epsilon}$
for all $t\geq 1$. Since $\widetilde{P}_{t}$ is a contraction, $\lVert\widetilde{P}_{t}\rVert_{\tau}\leq 1$ for all $t\geq 1$. As in the proof of Theorem 1.3, the series $\sum_{n=1}^{\infty}\frac{n^{k}}{2^{n/2}}$ converges, so taking $t\rightarrow+\infty$, we conclude that
$\lVert\widetilde{P}_{0}\rVert_{\tau}\leq\Gamma m^{k}\sqrt{\epsilon}$
for any integer $\Gamma\geq(C+25)\sum_{n=1}^{\infty}\frac{n^{k}}{2^{n/2}}$. ∎
Proposition 6.4.
Let $\mathcal{W}$ be the mapping from Lemma 3.11. For every $m\in\mathbb{N}$, let $\mathcal{W}_{m}=\big(\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\big)\cup\mathcal{W}(\widetilde{P}_{0})\cup\mathcal{R}_{m}$ be as in Proposition 5.4, and let
$\widetilde{\mathcal{W}}_{m}:=\{s\otimes 1:s\in\mathcal{W}_{m}\}\cup\{s\otimes 1,1\otimes s:s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)\}.$
There are positive integers $\widetilde{\Gamma},k^{\prime}$ such that if $m\notin\mathcal{L}$, $\varphi$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and $\varphi$ is $(\epsilon,\mathcal{X})$-synchronous, then $\lVert\widetilde{P}_{0}\otimes 1\rVert_{\varphi}\leq\widetilde{\Gamma}m^{k^{\prime}}\sqrt{\epsilon}$.
Proof.
Let $\Gamma,k$ be integers from Proposition 6.3, and let $D$ be the integer from Theorem 4.1, part (d). Suppose $m\not\in\mathcal{L}$, $\epsilon\geq 0$, and that $\varphi$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state on $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ which is $(\epsilon,\mathcal{X})$-synchronous. Since $\varphi(s^{*}s\otimes 1)\leq\epsilon$ and $\varphi(1\otimes s^{*}s)\leq\epsilon$ for all $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$, and $\{s\otimes 1:s\in\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\}\subseteq\widetilde{\mathcal{W}}_{m}$, by Lemma 3.12, there is a $(25\epsilon,\mathcal{X})$-synchronous state $\widehat{\varphi}$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that
$\big|\lVert r\otimes 1\rVert_{\widehat{\varphi}}-\lVert r\otimes 1\rVert_{\varphi}\big|\leq 2\lVert r\rVert_{1,1}\sqrt{\epsilon}\leq 2Dm\sqrt{\epsilon}$
for all $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$. Since $\{r\otimes 1:r\in\mathcal{R}_{m}\}\subseteq\widetilde{\mathcal{W}}_{m}$, for every $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$, $\lVert r\otimes 1\rVert_{\varphi}\leq\sqrt{\epsilon}$, and hence $\lVert r\otimes 1\rVert_{\widehat{\varphi}}\leq(2Dm+1)\sqrt{\epsilon}$. Rounding $2Dm+1$ up to $2Dm+3$ so that $(2Dm+3)^{2}\geq 25$, we conclude that $\widehat{\varphi}$ is a $((2Dm+3)^{2}\epsilon,\mathcal{X})$-synchronous state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ and $\widehat{\varphi}\circ\iota$ is a $((2Dm+3)^{2}\epsilon,\mathcal{R}_{m})$-state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By Proposition 6.3, $\lVert\widetilde{P}_{0}\otimes 1\rVert_{\widehat{\varphi}}\leq\Gamma m^{k}\sqrt{(2Dm+3)^{2}\epsilon}=(2Dm+3)\Gamma m^{k}\sqrt{\epsilon}$. Again by Lemma 3.12, since $\{s\otimes 1:s\in\mathcal{W}(\widetilde{P}_{0})\}\subseteq\widetilde{\mathcal{W}}_{m}$, we have
$\big|\lVert\widetilde{P}_{0}\otimes 1\rVert_{\widehat{\varphi}}-\lVert\widetilde{P}_{0}\otimes 1\rVert_{\varphi}\big|\leq 2\lVert\widetilde{P}_{0}\rVert_{1,1}\sqrt{\epsilon}\leq 3\sqrt{\epsilon}.$
It follows that
$\lVert\widetilde{P}_{0}\otimes 1\rVert_{\varphi}\leq\left((2Dm+3)\Gamma m^{k}+3\right)\sqrt{\epsilon}\leq 8D\Gamma m^{k+1}\sqrt{\epsilon}.$
Taking $\widetilde{\Gamma}=8D\Gamma$ and $k^{\prime}=k+1$ completes the proof. ∎
Before proceeding to the proof of Theorem 6.1, we need a few more lemmas. The first is a standard fact about extending tracial states to (not necessarily tracial) states on tensor products.
Lemma 6.5.
If $\tau$ is a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$, then there is a $(0,\mathcal{X})$-synchronous state $\varphi$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ with $\varphi\circ\iota=\tau$.
Proof.
Let $(\pi,v)$ be the GNS representation of $\tau$, and let $\mathcal{C}=\overline{\pi(\mathcal{A})}$ be the closure of $\pi(\mathcal{A})$ in the operator norm, where $\mathcal{A}=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$. By [BO08, Proposition 6.1.2], the tracial state defined by $v$ on $\mathcal{C}$ extends to a state $\varphi^{\prime}:\mathcal{C}\otimes\mathcal{C}^{op}\to\mathbb{C}:a\otimes b\mapsto\langle v,abv\rangle$. Note that the isomorphism $\omega_{\mathcal{X}}:\mathbb{C}^{*}\langle\mathcal{X}\rangle\to\mathbb{C}^{*}\langle\mathcal{X}\rangle^{op}$ defined in Section 3 descends to an isomorphism $\mathcal{A}\to\mathcal{A}^{op}$, and that $\pi$ defines a *-homomorphism $\pi^{op}:\mathcal{A}^{op}\to\mathcal{C}^{op}$. Let $\varphi=\varphi^{\prime}\circ(\pi\otimes(\pi^{op}\circ\omega_{\mathcal{X}}))$. Then $\varphi$ is a state on $\mathcal{A}\otimes\mathcal{A}$ with $\varphi(a\otimes b)=\langle v,\pi(a\omega_{\mathcal{X}}(b))v\rangle=\tau(a\cdot\omega_{\mathcal{X}}(b))$. Since
$\lVert x\otimes 1-1\otimes x\rVert_{\varphi}^{2}=\varphi(x^{*}x\otimes 1-x\otimes x^{*}-x^{*}\otimes x+1\otimes x^{*}x)=\tau(x^{2}-x^{2}-x^{2}+x^{2})=0$
for all $x\in\mathcal{X}$, $\varphi$ is $(0,\mathcal{X})$-synchronous. ∎
The remaining two lemmas are used to get from $\mathbb{Q}\mathcal{F}_{N}$ and $\mathbb{Q}\mathbb{Z}_{2}^{*N}$ for large $N$ down to $\mathbb{Q}\mathbb{Z}_{3}^{*2}$ and $\mathbb{Q}\mathbb{Z}_{2}^{*3}$. Both lemmas will be familiar to experts.
Lemma 6.6.
If $H$ be a subgroup of a group $G$, then any state on $\mathbb{C}H$ extends to a state on $\mathbb{C}G$.
Proof.
Let $\mathbb{C}H_{+}$ and $\mathbb{C}G_{+}$ be the convex cones of sums of squares in $\mathbb{C}H$ and $\mathbb{C}G$ respectively, and let $r:\mathbb{C}G\to\mathbb{C}H$ be the linear function with $r(g)=g$ if $g\in H$ and $r(g)=0$ if $g\not\in H$. By [SS11, Proposition 4], $r(\mathbb{C}G_{+})\subseteq\mathbb{C}H_{+}$. Consequently, if $a\in\mathbb{C}G_{+}\cap\mathbb{C}H$, then $a=r(a)\in\mathbb{C}H_{+}$, so $\mathbb{C}G_{+}\cap\mathbb{C}H=\mathbb{C}H_{+}$. If $f$ is a state on $\mathbb{C}H$, then $f$ is nonnegative on $\mathbb{C}H_{+}$. Since $\mathbb{C}G$ is archimedean, the M. Riesz extension theorem implies that $f$ extends to a real linear functional on $\text{span}_{\mathbb{R}}\mathbb{C}G_{+}$ which is nonnegative on $\mathbb{C}G_{+}$. From there, we can extend further to get a linear functional $\widetilde{f}$ on $\mathbb{C}G$ such that $\widetilde{f}(a^{*})=\overline{\widetilde{f}(a)}$ for all $a\in\mathbb{C}G$, $\widetilde{f}|_{\mathbb{C}H}=f$, and $\widetilde{f}$ is nonnegative on $\mathbb{C}G_{+}$. Since $\widetilde{f}(1)=f(1)=1$, the archimedean condition implies that $\widetilde{f}$ is a state. ∎
Another way to prove this lemma is to observe that $\mathbb{C}H$ and $\mathbb{C}G$ are subalgebras of the full group $C^{*}$-algebras $C^{*}(H)$ and $C^{*}(G)$. By definition of the full group $C^{*}$-algebra, any state on $\mathbb{C}H$ is the restriction of a state on $C^{*}(H)$, and since $C^{*}(H)$ is a subalgebra of $C^{*}(G)$, any state on $C^{*}(H)$ extends to a state on $C^{*}(G)$.
Lemma 6.7.
Suppose $N\geq 1$ and $G$ is either $\mathcal{F}_{K}$ or $\mathbb{Z}_{3}^{*K}$ with $K\geq 2$, or $\mathbb{Z}_{2}^{*K}$ with $K\geq 3$. Then there is an injective group homomorphism $\mu:\mathcal{F}_{N}\to G$ and a computable *-homomorphism $\widetilde{\mu}:\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle\to\mathbb{Q}^{*}\langle x_{1},\ldots,x_{K}\rangle$ such that the diagram
[…]
commutes, where $q_{1}$ and $q_{2}$ are the quotient homomorphisms sending $x_{j}$ to the $j$th generator of $\mathcal{F}_{N}$ and $G$ respectively.
Proof.
Let $x_{1},\ldots,x_{N}$ denote the generators of $\mathcal{F}_{N}$, $y_{1},y_{2}$ denote the generators of $\mathcal{F}_{2}$, $w_{1},w_{2}$ denote the generators of $\mathbb{Z}_{3}^{*2}$, and $z_{1},z_{2},z_{3}$ denote the generators of $\mathbb{Z}_{2}^{*3}$. It is well-known that the homomorphisms
$\gamma_{1}:\mathcal{F}_{N}\to\mathcal{F}_{2}:x_{i}\mapsto y_{1}^{i}y_{2}y_{1}^{-i},1\leq i\leq N,$
$\gamma_{2}:\mathcal{F}_{2}\to\mathbb{Z}_{3}^{*2}:y_{1}\mapsto w_{1}w_{2}^{-1},y_{2}\mapsto w_{1}^{-1}w_{2},\text{ and}$
$\gamma_{3}:\mathcal{F}_{2}\to\mathbb{Z}_{2}^{*3}:y_{i}\mapsto z_{1}^{1-i}z_{2}z_{3}z_{1}^{1-i},i=1,2$
are injective. If $G=\mathcal{F}_{K}$ with $K\geq 2$, then we can take $\mu=\gamma_{1}$, where we think of $\mathcal{F}_{2}$ as a subgroup of $\mathcal{F}_{K}$ by mapping $y_{1}$ and $y_{2}$ to the first two generators, and let $\widetilde{\mu}$ be the lift of $\mu$ sending $x_{i}\mapsto x_{1}^{i}x_{2}(x_{1}^{*})^{i}$. If $G=\mathbb{Z}_{3}^{*K}$ with $K\geq 2$, we can take $\mu=\gamma_{2}\circ\gamma_{1}$, where again we think of $\mathbb{Z}_{3}^{*2}$ as a subgroup of $\mathbb{Z}_{3}^{*K}$ by mapping $w_{1}$ and $w_{2}$ to the first two generators, and let $\widetilde{\mu}$ be the lift of $\mu$ sending $x_{i}\mapsto(w_{1}w_{2}^{*})^{i}w_{1}^{*}w_{2}(w_{2}w_{1}^{*})^{i}$. Finally if $G=\mathbb{Z}_{2}^{*K}$ with $K\geq 3$, we can take $\mu=\gamma_{3}\circ\gamma_{1}$, where we think of $\mathbb{Z}_{2}^{*3}$ as a subgroup of $\mathbb{Z}_{2}^{*K}$, and let $\widetilde{\mu}$ be the lift of $\mu$ sending $x_{i}\mapsto(x_{2}x_{3})^{i}(x_{1}x_{2}x_{3}x_{1})(x_{3}^{*}x_{2}^{*})^{i}$. In all three cases, we see from the explicit formula that $\widetilde{\mu}$ is computable. ∎
Proof of Theorem 6.1.
Set $\mathcal{P}_{N}:=\mathbb{Q}^{*}\langle x_{1},\ldots,x_{N}\rangle$ for $N\in\mathbb{N}$. Let $P_{1}(N)$ be the promise problem from the statement of the theorem with $\mathcal{P}=\mathcal{P}_{N}$ and $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{2}^{*N}$, and let $P_{2}(N)$ be the promise problem with $\mathcal{P}=\mathcal{P}_{N}$ and $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{3}^{*N}$. Although it isn’t mentioned in the theorem statement, let $P_{3}(N)$ be the promise problem with $\mathcal{P}=\mathcal{P}_{N}$ and $\mathcal{C}=\mathbb{Q}\mathcal{F}_{N}$. As in the proof of Theorem 5.1, let $n_{0}=\lvert\mathcal{X}\rvert$. We start by showing that $P_{1}(N)$ is coRE-hard for all $N\geq n_{0}$.
For this, suppose $N\geq n_{0}$, and once again pick an ordering on $\mathcal{X}$ and identify $\mathbb{Q}^{*}\langle\mathcal{X}\rangle$, $\mathbb{C}^{*}\langle\mathcal{X}\rangle$, and $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ with subalgebras of $\mathcal{P}_{N}$, $\mathbb{C}^{*}\langle x_{1},\ldots,x_{N}\rangle$, and $\mathbb{C}\mathbb{Z}_{2}^{*N}$ respectively. Let $\pi:\mathcal{P}_{N}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ be the quotient map, and let $q=q_{2}\circ q_{1}$, where $q_{1}:\mathbb{Q}\mathbb{Z}_{2}^{*N}\to\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}$ is the homomorphism sending $x_{i}\mapsto x_{i}$ for $1\leq i\leq n_{0}$, and $q_{2}:\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\to\mathcal{A}_{\mathcal{L}}(m)$ is the quotient map. Let $\widetilde{\Gamma},k^{\prime}$, and $\widetilde{W}_{m}$ be as in Proposition 6.4. For every $m\in\mathbb{N}$, define a *-polynomial
$\beta(m)=\sum_{r\in\widetilde{\mathcal{W}}_{m}\cup\widetilde{\mathcal{W}}_{m}^{*}}r^{*}r+\sum_{x\in\mathcal{X}}(x\otimes 1-1\otimes x)^{*}(x\otimes 1-1\otimes x)-\tfrac{1}{\widetilde{\Gamma}^{2}m^{2k^{\prime}}}\widetilde{P}_{0}^{*}\widetilde{P}_{0}\otimes 1$
in $\mathbb{Q}^{*}\langle\mathcal{X}\rangle\otimes_{\mathbb{Q}}\mathbb{Q}^{*}\langle\mathcal{X}\rangle\subseteq\mathcal{P}_{N}\otimes_{\mathbb{Q}}\mathcal{P}_{N}$. Then $\beta(m)$ is self-adjoint and the map $\beta:\mathbb{N}\to\mathcal{P}_{N}\otimes_{\mathbb{Q}}\mathcal{P}_{N}$ is computable by Lemmas 3.11, 4.1 and 6.4.
By Lemma 3.11, $(\pi\otimes\pi)(t\otimes 1)=(\pi\otimes\pi)(s\otimes 1)=(\pi\otimes\pi)(1\otimes s)=0$ for all $t\in\big(\bigcup_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}\mathcal{W}(r)\big)\cup\mathcal{W}(\widetilde{P}_{0})$ and $s\in\bigcup_{x\in\mathcal{X}}\mathcal{W}(x)$. Hence
$(\pi\otimes\pi)(\beta(m))=\sum_{r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}}r^{*}r\otimes 1+\sum_{x\in\mathcal{X}}(x\otimes 1-1\otimes x)^{2}-\tfrac{1}{\widetilde{\Gamma}^{2}m^{2k^{\prime}}}\widetilde{P}_{0}^{2}\otimes 1$
in $\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes_{\mathbb{Q}}\mathbb{Q}\mathbb{Z}_{2}^{*\mathcal{X}}\subseteq\mathbb{Q}\mathbb{Z}_{2}^{*N}\otimes_{\mathbb{Q}}\mathbb{Q}\mathbb{Z}_{2}^{*N}$. If $m\in\mathcal{L}$, then by Theorem 4.1, part (c), there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(\widetilde{P}_{0}^{2})>0$. Since $\tau\circ q_{2}$ is a tracial state on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ and $q_{2}(r)=0$ for all $r\in\mathcal{R}_{m}$, Lemma 6.5 implies that there is a $(0,\mathcal{X})$-synchronous state $\varphi$ on $\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}\otimes\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}}$ such that $\varphi(\widetilde{P}_{0}^{2}\otimes 1)=\tau(\widetilde{P}_{0}^{2})>0$ and $\varphi(r^{*}r\otimes 1)=\tau(q_{2}(r^{*}r))=0$ for all $r\in\mathcal{R}_{m}\cup\mathcal{R}_{m}^{*}$. Then $\widetilde{\varphi}:=\varphi\circ(q_{1}\otimes q_{1})$ is a state on $\mathbb{C}\mathbb{Z}_{2}^{*N}\otimes\mathbb{C}\mathbb{Z}_{2}^{*N}$ such that $\widetilde{\varphi}\left((\pi\otimes\pi)(\beta(m))\right)=-\tfrac{1}{\widetilde{\Gamma}^{2}m^{2k^{\prime}}}\tau(\widetilde{P}_{0}^{2})<0$. We conclude that $(\pi\otimes\pi)(\beta(m))$ is not positive.
Now suppose $m\notin\mathcal{L}$. Let $\varphi$ be a state on $(\mathcal{P}_{N})_{\mathbb{C}}\otimes_{\mathbb{C}}(\mathcal{P}_{N})_{\mathbb{C}}$, and let
$\epsilon:=\varphi\left(\sum_{r\in\widetilde{\mathcal{W}}_{m}\cup\widetilde{\mathcal{W}}_{m}^{*}}r^{*}r+\sum_{x\in\mathcal{X}}(x\otimes 1-1\otimes x)^{*}(x\otimes 1-1\otimes x)\right).$
The restriction of $\varphi$ to $\mathbb{C}^{*}\langle\mathcal{X}\rangle\otimes\mathbb{C}^{*}\langle\mathcal{X}\rangle$ is an $(\epsilon,\widetilde{\mathcal{W}}_{m})$-state which is $(\epsilon,\mathcal{X})$-synchronous, so by Proposition 6.4, $\varphi(\widetilde{P}_{0}^{*}\widetilde{P}_{0}\otimes 1)=\lVert\widetilde{P}_{0}\otimes 1\rVert^{2}_{\varphi}\leq\widetilde{\Gamma}^{2}m^{2k^{\prime}}\epsilon$. Hence $\varphi(\beta(m))\geq 0$. We conclude that the complement $\overline{\mathcal{L}}$ of $\mathcal{L}$ reduces to $P_{1}(N)$. Taking $\mathcal{L}$ to be RE-hard implies that $P_{1}(N)$ is coRE-hard for $N\geq n_{0}$.
Next, the quotient map $\pi^{(N)}:\mathcal{P}_{N}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$ factors as $\pi^{(N)}=\pi_{2}^{(N)}\circ\pi_{1}^{(N)}$, where $\pi_{1}^{(N)}$ is the quotient map $\mathcal{P}_{N}\to\mathbb{Q}\mathcal{F}_{N}$, and $\pi_{2}^{(N)}$ is the quotient map $\mathbb{Q}\mathcal{F}_{N}\to\mathbb{Q}\mathbb{Z}_{2}^{*N}$. If $(\pi\otimes\pi)(\alpha)$ is not positive in $\mathbb{C}\mathbb{Z}_{2}^{*N}\otimes\mathbb{C}\mathbb{Z}_{2}^{*N}$ for $\alpha\in\mathcal{P}_{N}\otimes\mathcal{P}_{N}$, then $(\pi_{1}^{(N)}\otimes\pi_{1}^{(N)})(\alpha)$ is not positive in $\mathbb{C}\mathcal{F}_{N}\otimes\mathbb{C}\mathcal{F}_{N}$. Thus $P_{1}(N)$ reduces to $P_{3}(N)$, and $P_{3}(N)$ is coRE-hard for $N\geq n_{0}$.
By Lemma 6.7, if $N\geq 1$ and $K\geq 2$ then there is an injective homomorphism $\mu:\mathcal{F}_{N}\to\mathbb{Z}_{3}^{*K}$ and a computable homomorphism $\widetilde{\mu}:\mathcal{P}_{N}\to\mathcal{P}_{K}$ such that $\pi_{3}^{(K)}\circ\widetilde{\mu}=\mu\circ\pi_{1}^{(N)}$, where $\pi_{3}^{(K)}$ is the quotient map $\mathcal{P}_{K}\to\mathbb{Q}\mathbb{Z}_{3}^{*K}$. If $\alpha\in\mathcal{P}_{N}\otimes\mathcal{P}_{N}$ is positive in $(\mathcal{P}_{N})_{\mathbb{C}}\otimes(\mathcal{P}_{N})_{\mathbb{C}}$, then $(\widetilde{\mu}\otimes\widetilde{\mu})(\alpha)$ is positive in $(\mathcal{P}_{K})_{\mathbb{C}}\otimes(\mathcal{P}_{K})_{\mathbb{C}}$. If $(\pi_{1}^{(N)}\otimes\pi_{1}^{(N)})(\alpha)$ is not positive in $\mathbb{C}\mathcal{F}_{N}\otimes\mathbb{C}\mathcal{F}_{N}$, then there is a state $f$ on $\mathbb{C}\mathcal{F}_{N}\times\mathcal{F}_{N}$ such that $f(\pi_{1}^{(N)}\otimes\pi_{1}^{(N)})(\alpha))<0$. If we identify $\mathbb{C}\mathcal{F}_{N}\otimes\mathbb{C}\mathcal{F}_{N}$ with $\mathbb{C}\mathcal{F}_{N}\times\mathcal{F}_{N}$ and $\mathbb{C}\mathbb{Z}_{3}^{*K}\otimes\mathbb{C}\mathbb{Z}_{3}^{*K}$ with $\mathbb{C}\mathbb{Z}_{3}^{*K}\times\mathbb{Z}_{3}^{*K}$, then the homomorphism $\mu\otimes\mu$ is the algebra homomorphism coming from the injective group homomorphism $\mu\times\mu:\mathcal{F}_{N}\times\mathcal{F}_{N}\to\mathbb{Z}_{3}^{*K}\times\mathbb{Z}_{3}^{*K}$. By Lemma 6.6, the state $f$ extends to a state $\widetilde{f}$ on $\mathbb{C}\mathbb{Z}_{3}^{*K}\times\mathbb{Z}_{3}^{*K}$, and
$\widetilde{f}((\mu\times\mu)(\pi_{1}^{(N)}\otimes\pi_{1}^{(N)})(\alpha))=f((\pi_{1}^{(N)}\otimes\pi_{1}^{(N)})(\alpha))<0.$
Hence $(\pi_{3}^{(K)}\otimes\pi_{3}^{(K)})(\widetilde{\mu}\otimes\widetilde{\mu})(\alpha)=(\mu\otimes\mu)(\pi_{1}^{(N)}\otimes\pi_{1}^{(N)})(\alpha)$ is not positive. Since $\widetilde{\mu}\otimes\widetilde{\mu}$ is computable, $P_{3}(N)$ reduces to $P_{2}(K)$ for any $N\geq 1$ and $K\geq 2$. It follows that $P_{2}(K)$ is coRE-hard for all $K\geq 2$. Similarly, $P_{3}(N)$ reduces to $P_{1}(K)$ for all $N\geq 1$ and $K\geq 3$, so $P_{1}(K)$ is coRE-hard for all $K\geq 3$. ∎
The following example shows that Theorem 6.1 is not true for $\mathcal{C}=\mathbb{Q}\mathbb{Z}_{2}^{*2}$.
Example 6.8.
Let $G=\mathbb{Z}_{2}^{*2}$ be the infinite dihedral group. This group is amenable and linear. Hence the product group $G\times G$ is also amenable and linear. By a theorem of Bekka, $\mathbb{C}G\times G$ is residually finite-dimensional [BEK99], in the sense that an element $\alpha\in\mathbb{C}G\otimes\mathbb{C}G=\mathbb{C}G\times G$ is positive if and only if $\phi(\alpha)$ is positive for all finite-dimensional representations $\phi$. Every finite-dimensional representation of $\mathbb{C}G\times G$ decomposes into irreducible representations, so $\alpha$ is positive if and only if $\phi(\alpha)$ is positive for all irreducible finite-dimensional representations. The irreducible representations of $G$ have dimension $1$ and $2$, so the irreducible representations of $G\times G$ have dimension $1$, 22, and $4$. Thus $\alpha\in\mathbb{C}G\times G$ is positive if and only if $\phi(\alpha)$ is positive for all representations of dimension $4$. The question of whether $\phi(\alpha)$ is not positive for all representations of dimension $4$ lies in the existential theory of the reals, and thus can be solved in $\operatorname{PSPACE}$ [CAN88]. In particular, the question of whether $\alpha\in\mathbb{C}G\times G$ is positive is decidable. We do not know whether $\mathbb{C}G\times G$ is archimedean closed.